### v2

# Chains of model structures arising from modules of finite Gorenstein dimension

Abstract.
For any integer $n\geq 0$ and any ring $R$,  $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ proves to be a complete hereditary cotorsion pair in $R$-Mod, where $\mathcal{PGF}$ is the class of PGF modules, introduced by J. Šaroch and J. Št́ovíček, and  $\mathcal{PGF}_{n}$ is the class of $R$-modules of PGF dimension $\leq n$. For any Artin algebra $R$,  $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp})$ proves to be a complete and hereditary cotorsion pair in $R$-Mod, where $\mathcal{GP}_{n}$ is the class of modules of Gorenstein projective dimension $\leq n$. These cotorsion pairs induce two chains of hereditary Hovey triples  $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ and  $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{GP}^{\perp})$, and the corresponding homotopy categories in the same chain are the same. It is observed that some complete cotorsion pairs in $R$-Mod can induce complete cotorsion pairs in some special extension closed subcategories of $R$-Mod. Then corresponding results in exact categories $\mathcal{PGF}_{n}$,  $\mathcal{GP}_{n}$,  $\mathcal{GF}_{n}$,  $\mathcal{PGF}^{<\infty}$,  $\mathcal{GP}^{<\infty}$ and $\mathcal{GF}^{<\infty}$, are also obtained. As a byproduct, $\mathcal{PGF}=\mathcal{GP}$ for a ring $R$ if and only if $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$ for some $n$.
Keywords: projectively coresolved Gorenstein flat (PGF) module; module of finite PGF (Gorenstein projective, Gorenstein flat) dimension; (complete, hereditary) cotorsion pair; abelian (exact) model structure; weakly idempotent exact category; finitistic dimension
2020 Mathematics Subject Classification. Primary 16E30, 18N40; Secondary 16E10, 16E65, 16G50
$*{}^{*}$ Corresponding author
[…]
Supported by National Natural Science Foundation of China, Grant No. 12131015, 12271333, and Natural Science Foundation of Shanghai under Grant No. 23ZR1435100.
Introduction
Gorenstein projective modules, introduced by M. Auslander and M. Bridger [1], and E. Enochs and O. M. G. Jenda [10], have got deep interests and wide applications in mathematics. A recent milestone is the introduction of projectively coresolved Gorenstein flat (PGF, for short) modules over an arbitrary ring $R$, by J. Šaroch and J. Št́ovíček [28].
This kind of PGF modules enjoys pleasant properties. It is not clear whether a Gorenstein projective module is Gorenstein flat. But PGF modules provide a common refinement of Gorenstein projective modules and Gorenstein flat modules: by definition they are Gorenstein flat, and they also prove to be Gorenstein projective ([28, Theorem 4.4]). The full subcategory $\mathcal{PGF}$ of $R$-Mod consisting of the PGF modules is a Frobenius category with $\mathcal{P}$, the full subcategory of projective $R$-modules, as the class of projective-injective objects, and the stable triangulated category $\mathcal{PGF}/\mathcal{P}$ is the homotopy category of infinite different abelian or exact model structures, in the sense of D. Quillen [24], M. Hovey [19], and J. Gillespie [14].
What interesting and important are, all the three classes $\mathcal{PGF}$, $\mathcal{GP}$, and $\mathcal{GF}$, are the left parts of cotorsion pairs in $R$-Mod, where $\mathcal{GP}$ (respectively, $\mathcal{GF}$) is the full subcategory of $R$-Mod consisting of the Gorenstein projective (respectively, Gorenstein flat) modules. A. Beligiannis and I. Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A$-Mod, if $A$ is an Artin algebra. Šaroch and Št́ovíček [28, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$. By a result of A. Iacob [20, Proposition 9] (see also [22, Proposition 4.11]), if $R$ is a Gorenstein ring (i.e., $R$ is a left and right noetherian ring and the injective dimensions of ${}_{R}R$ and $R_{R}$ are finite), then $\mathcal{PGF}=\mathcal{GP}$. Thus, in this case, the cotorsion pair $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ reads as $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, which has been given by Hovey [19, Theorem 8.3]. By the Hovey correspondence, all these three kinds of cotorsion pairs induce abelian model structures in $R$-Mod.
Various homological dimensions provide more possibilities to obtain cotorsion pairs. For examples, for each non-negative integer $n$, both
$(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})\ \ \ \ \mbox{ and}\ \ \ \ (\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$
are complete and hereditary cotorsion pairs in $R$-Mod (see [11, Theorem 7.4.6], and [23, Theorem 3.4(2)], respectively), where $\mathcal{P}_{n}$ (respectively, $\mathcal{F}_{n})$ is the class of $R$-modules of projective (respectively, flat) dimension $\leq n$.
The first aim of this paper is to show that these phenomena are quite common. For each non-negative integer $n$, we prove that
$(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
is a complete and hereditary cotorsion pair in $R$-Mod for any ring $R$, where $\mathcal{PGF}_{n}$ is the class of $R$-modules of PGF dimension $\leq n$ (see Corollary 2.6); and that
$(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{P}^{<\infty})$
is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra, where $\mathcal{GP}_{n}$ is the class of $R$-modules of Gorenstein projective dimension $\leq n$ (see Theorem 3.4). R. EI Maaouy [22, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
$(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
for any ring $R$, where $\mathcal{GF}_{n}$ is the class of $R$-modules of Gorenstein flat dimension $\leq n$. These three chains of cotorsion pairs induce three chains of hereditary abelian model structures on $R$-Mod:
$(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp}),\ \ \ \ \ (\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{P}^{<\infty}),\ \ \ \ \ (\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp}).$
which are not trivial (i.e., not every module is a trivial object in the model structure), and not projective (i.e., not every module is a fibrant object) in general. (We remind that, in this paper, a Hovey triple is written in the order  $(\mathcal{C},\mathcal{F},\mathcal{W})$, rather than  $(\mathcal{C},\mathcal{W},\mathcal{F})$, where $\mathcal{W}$ is the class of trivial objects.) Since all the hereditary abelian model structures in each chain have the same class of trivial objects, the homotopy categories in the same chain are triangle equivalent to each other. Thus the homotopy categories are the stable categories $\mathcal{PGF}/\mathcal{P}$ and $\mathcal{GP}/\mathcal{P}$.
One may further ask whether $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$, and $\mathcal{GF}^{<\infty}$, can be the left parts of complete and hereditary cotorsion pairs in $R$-Mod, where $\mathcal{PGF}^{<\infty}$ (respectively, $\mathcal{GP}^{<\infty},\ \mathcal{GF}^{<\infty}$) is the class of $R$-modules of finite PGF (respectively, the Gorenstein projective, Gorenstein flat) dimension. This leads to an investigation on various finitistic dimensions. See Proposition 4.1 and Corollary 4.2.
An overall landscape of this question can be viewed in the following table.
Table 1:  Complete and hereditary cotorsion pairs
               and the induced Hovey triples in $R$-Mod
$n=0$
$n$
<absent<$\infty$< ∞
$\mathcal{P}$
$(\mathcal{P},\ R\mbox{-}{\rm Mod})$
$(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$
Enochs, Jenda [11, Theorem 7.4.6]
If ${\rm Fpd}<\infty$, then
 $(\mathcal{P}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp})$
$\mathcal{FL}$
$=\mathcal{F}_{0}$
$(\mathcal{FL},\ \mathcal{EC})$
Enochs, Jenda [11, Proposition 7.4.3]
$(\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$
Mao, Ding [23, Theorem 3.4(2)]
If ${\rm Fpd}<\infty$, then
 $(\mathcal{F}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp})$
$\mathcal{PGF}$
$(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$
$(\mathcal{PGF},\ R\mbox{-}{\rm Mod},\ \mathcal{PGF}^{\perp})$
Šaroch, Št́ovíček [28, Theorem 4.9]
$(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
$(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$
Corollary 2.6
If ${\rm Fpd}<\infty$, then
 $(\mathcal{PGF}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})$
$(\mathcal{PGF}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})$
Corollary 4.2
$\mathcal{GP}$
For Artin algebras
$(\mathcal{GP},\ \mathcal{GP}^{\perp})$
$(\mathcal{GP},\ R\mbox{-}{\rm Mod},\ \mathcal{GP}^{\perp})$
Beligiannis, Reiten
[5, X, Theorem 2.4(iv)]
For Artin algebras
$(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp})$
$(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{GP}^{\perp})$
Theorem 3.4
For Artin algebras
If ${\rm Fpd}<\infty$, then
$(\mathcal{GP}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp}\cap\mathcal{GP}^{\perp})$
$(\mathcal{GP}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp},\ \mathcal{GP}^{\perp})$
Corollary 4.2
$\mathcal{GF}$
$(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$
$(\mathcal{GF},\ \mathcal{EC},\ \mathcal{PGF}^{\perp})$
Šaroch, Št́ovíček
[28, Theorem 4.11, Corollary 4.12]
$(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
$(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$
Maaouy [22, Theorem A]
If ${\rm Fpd}<\infty$, then
$(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})$
$(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})$
Corollary 4.2
In order to study homology in non abelian categories, Quillen [26] introduces exact categories. Any full subcategory of an abelian category which are closed under extensions and direct summands is a weakly idempotent complete exact category. By Gillespie [14], there is also the Hovey correspondence between the Hovey triples and the exact model structures, on any weakly idempotent complete exact category (see also [27]).
The second aim of this paper is to look for complete and hereditary cotorsion pairs and exact model structures in weakly idempotent complete exact categories $\mathcal{PGF}_{n}$, $\mathcal{GP}_{n}$, $\mathcal{GF}_{n}$, and $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$ and $\mathcal{GF}^{<\infty}$, for any non-negative integer $n$. This mainly comes from the following observation (see Theorem 5.1):
If $\mathcal{B}$ is a full subcategory of abelian category $\mathcal{A}$, which is closed under extensions and the kernels of epimorphisms, then any complete cotorsion pair $(\mathcal{X},\ \mathcal{Y})$ in $\mathcal{A}$ with $\mathcal{X}\subseteq\mathcal{B}$ induces complete cotorsion pair $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ in exact category $\mathcal{B}.$ Moreover, if $(\mathcal{X},\ \mathcal{Y})$ is hereditary in $\mathcal{A}$, then so is $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ in $\mathcal{B}$.
This observation provides new complete cotorsion pairs and exact model structures in exact categories. For example, for each non-negative integer $m$, one has two chains of complete and hereditary cotorsion pairs $(\mathcal{X},\ \mathcal{Y})$ in $R$-Mod with $\mathcal{X}\subseteq\mathcal{PGF}^{<\infty}:$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}),\ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}).$
Applying Theorem 5.1 one gets two chains of complete and hereditary cotorsion pairs in $\mathcal{PGF}^{<\infty}$:
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{<\infty}),\ \ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}^{<\infty});$
and one chain of hereditary Hovey triples $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{<\infty},\ \mathcal{P}^{<\infty})$ in $\mathcal{PGF}^{<\infty}$. See Corollary 5.4. When $m=0$, this is [8, Theorem 4.1] by G. Dalezios and I. Emmanouil.
Let us see the case in $\mathcal{GP}_{n}$. For any non-negative integer $n$, one has two chains of complete and hereditary cotorsion pairs $(\mathcal{X},\ \mathcal{Y})$ in $R$-Mod with $\mathcal{X}\subseteq\mathcal{GP}_{n}$:
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}),\ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}),$
where $0\leq m\leq n$. Applying Theorem 5.1 one gets two chains of complete and hereditary cotorsion pairs in $\mathcal{GP}_{n}:$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n}),\ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n});$
and one chain of hereditary Hovey triples $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ in $\mathcal{GF}_{n}$. See Corollary 6.1. In this way one gets double chains of complete and hereditary cotorsion pairs and hereditary Hovey triples, parameterized by $m$ and $n$. When $m=0=n$, this is G. Dalezios and I. Emmanouil [8, Theorem 4.3]. All these exact model structures are not trivial, and not projective in general, with the same homotopy category $\mathcal{PGF}/\mathcal{P}$.
More interesting, $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is a complete and hereditary cotorsion pair in $\mathcal{GP}_{n}$; and $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{P}_{n})$ is a hereditary Hovey triple in $\mathcal{GP}_{n}$, for any integers $m$ and $n$ with $0\leq m\leq n$. See Theorem 6.3. What surprising is, Theorem 6.3 is for an arbitrary ring $R$, not only for Artin algebras. Thus, Theorem 6.3 is not an application of Theorem 5.1 and Theorem 3.4 (which is for Artin algebras), and its proof is different from the one of Corollary 6.1.
As a byproduct, for a ring $R$, $\mathcal{PGF}=\mathcal{GP}$ if and only if $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$ for some non-negative integer $n$; if and only if $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$ for any non-negative integer $n$.
An overall landscape can be viewed in the following table.
Table 2:  Complete and hereditary cotorsion pairs and
              the induced Hovey triples in exact categories
$n=0$
$n$
$0\leq m\leq n$
<absent<$\infty$< ∞
$m$
$\mathcal{P}$
In $\mathcal{P}$
$(\mathcal{P},\ \mathcal{P})$
In $\mathcal{P}_{n}$
$(\mathcal{P}_{m},\ \mathcal{P}^{\perp}_{m}\cap\mathcal{P}_{n})$
In $\mathcal{P}^{<\infty}$
$(\mathcal{P}_{m},\ \mathcal{P}^{\perp}_{m}\cap\mathcal{P}^{<\infty})$
$\mathcal{FL}=\mathcal{F}_{0}$
In $\mathcal{FL}$
$(\mathcal{FL},\ \mathcal{EC}\cap\mathcal{FL})$
In $\mathcal{F}_{n}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{F}_{n})$
$(\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{F}_{n})$
In $\mathcal{F}^{<\infty}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{F}^{<\infty})$
$(\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{F}^{<\infty})$
$\mathcal{PGF}$
In $\mathcal{PGF}$
$(\mathcal{PGF},\ \mathcal{P})$
$(\mathcal{PGF},\ \mathcal{PGF},\ \mathcal{P})$
a special case of Corollary 5.3
In $\mathcal{PGF}_{n}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}_{n})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}_{n},\ \mathcal{P}_{n})$
Corollary 5.3
In $\mathcal{PGF}^{<\infty}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{<\infty})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}^{<\infty})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{<\infty},\ \mathcal{P}^{<\infty})$
Corollary 5.4
$\mathcal{GP}$
In $\mathcal{GP}$
$(\mathcal{PGF},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP})$
$(\mathcal{GP},\ \mathcal{P})$
$(\mathcal{PGF},\ \mathcal{GP},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP})$
$(\mathcal{GP},\ \mathcal{GP},\ \mathcal{P})$
a special case of
Corollary 6.1 and Theorem 6.3
In $\mathcal{GP}_{n}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$
$(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$
$(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{P}_{n})$
Corollary 6.1 and Theorem 6.3
In $\mathcal{GP}^{<\infty}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{<\infty})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GP}^{<\infty})$
$(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}^{<\infty})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{<\infty},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}^{<\infty})$
$(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{<\infty},\ \mathcal{P}^{<\infty})$
Corollary 6.5
$\mathcal{GF}$
In $\mathcal{GF}$
$(\mathcal{FL},\ \mathcal{EC}\cap\mathcal{GF})$
$(\mathcal{GF},\ \mathcal{FL}\cap\mathcal{EC})$
$(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{GF},\ \mathcal{FL})$
a special case of Corollary 7.7
$(\mathcal{PGF},\ \mathcal{FL})$
$(\mathcal{PGF},\ \mathcal{GF},\ \mathcal{FL})$
Dalezios, Emmanouil
[8, Theorem 4.3]
In $\mathcal{GF}_{n}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}_{n})$
$(\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}_{n})$
$(\mathcal{PGF}_{m},\ \mathcal{F}_{n}\cap\mathcal{P}_{m}^{\perp})$
$(\mathcal{GF}_{m},\ \mathcal{F}_{n}\cap\mathcal{F}_{m}^{\perp})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}_{n},\ \mathcal{F}_{n})$
$(\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}_{n},\ \mathcal{F}_{n})$
Corollary 7.7
In $\mathcal{GF}^{<\infty}$
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}^{<\infty})$
$(\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}^{<\infty})$
$(\mathcal{PGF}_{m},\ \mathcal{F}^{<\infty}\cap\mathcal{P}_{m}^{\perp})$
$(\mathcal{GF}_{m},\ \mathcal{F}^{<\infty}\cap\mathcal{F}_{m}^{\perp})$
$(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}^{<\infty},\ \mathcal{F}^{<\infty})$
$(\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}^{<\infty},\ \mathcal{F}^{<\infty})$
Corollary 7.9
The paper is organized as follows.
1.  Preliminaries
2.  Abelian model structures arising from $\mathcal{PGF}_{n}$
3.  Abelian model structures arising from $\mathcal{GP}_{n}$
4.  Several finitistic dimensions
4.1  Equalities on finitistic dimensions
4.2  Abelian model structures arising from $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$, and $\mathcal{GF}^{<\infty}$
5.  Induced complete cotorsion pairs in exact categories with applications
5.1  Induced complete cotorsion pairs in exact categories
5.2  Application 1: Exact model structures on $\mathcal{PGF}_{n}$
5.3  Application 2: Exact model structures on $\mathcal{PGF}^{<\infty}$
6.  Exact model structures on $\mathcal{GP}_{n}$ and $\mathcal{GP}^{<\infty}$
6.1   Application 3: Exact model structures on $\mathcal{GP}_{n}$
6.2   A new chain of cotorsion pairs and exact model structures on $\mathcal{GP}_{n}$
6.3   A sufficient and necessary condition for $\mathcal{PGF}=\mathcal{GP}$
6.4  Exact model structures on $\mathcal{GP}^{<\infty}$
7.  Exact model structures on $\mathcal{GF}_{n}$ and $\mathcal{GF}^{<\infty}$
7.1  Application 4: Exact model structures on $\mathcal{GF}_{n}$
7.2  Application 5: Exact model structures on $\mathcal{GF}^{<\infty}$
1. Preliminaries
1.1. Modules of finite Gorenstein dimension
Throughout $R$ is a ring with identity, $R$-Mod the category of left $R$-modules, and $\mathcal{P}(R)$, or simply $\mathcal{P}$, the class of projective $R$-modules. Denote by $\mathcal{P}^{<\infty}$ (respectively, $\mathcal{F}^{<\infty}$) the class of modules of finite projective (respectively, flat) dimension. For a non-negative integer $n$, let $\mathcal{P}_{n}$ (respectively, $\mathcal{F}_{n}$) be the class of modules $M$ of projective dimension ${\rm pd}M\leq n$ (respectively, flat dimension ${\rm fd}M\leq n$).
An $R$-module $M$ is Gorenstein projective ([1]; [11]) if it is a syzygy of an exact sequence of projective modules which remains exact after applying ${\rm Hom}_{R}(-,P)$ for any projective $R$-module $P$. Let $\mathcal{GP}$ be the subcategory of Gorenstein projective $R$-modules. The Gorenstein projective dimension of a module $M$ is denoted by ${\rm Gpd}M$. Let $\mathcal{GP}^{<\infty}$ be the class of modules of finite Gorenstein projective dimension. For a non-negative integer $n$, let $\mathcal{GP}_{n}$ be the subcategory of $R$-modules $M$ with ${\rm Gpd}M\leq n$.
An $R$-module $M$ is Gorenstein flat ([11]) if it is a syzygy of an exact sequence of flat modules which remains exact after tensoring with any right injective $R$-module. It is open whether a Gorenstein projective module is Gorenstein flat. Denote by $\mathcal{GF}$ the class of Gorenstein flat $R$-modules. The Gorenstein flat dimension of module $M$ is denoted by ${\rm Gfd}M$ ([17]). Let $\mathcal{GF}^{<\infty}$ be the class of modules of finite Gorenstein flat dimension. For a non-negative integer $n$, let $\mathcal{GF}_{n}$ be the class of modules $M$ of ${\rm Gfd}M\leq n$.
Following [28], an $R$-module $M$ is projectively coresolved Gorenstein flat (PGF for short), if it is a syzygy of an exact sequence of projective modules which remains exact after tensoring with any right injective $R$-module. Denote by $\mathcal{PGF}$ the subcategory of PGF modules. By definition $\mathcal{PGF}\subseteq\mathcal{GF}.$ It is proved in [28, Theorem 4.4] that $\mathcal{PGF}\subseteq\mathcal{GP}.$ However, it is not clear whether a Gorenstein projective module is PGF. In fact, $\mathcal{PGF}=\mathcal{GP}$ if and only if $\mathcal{GP}\subseteq\mathcal{GF}$. Denoted by ${\rm PGFd}M$ the PGF dimension of module $M$ (see [8]). Denote by $\mathcal{PGF}^{<\infty}$ the class of modules of finite PGF dimension. For a non-negative integer $n$, let $\mathcal{PGF}_{n}$ be the subcategory of $R$-modules $M$ with ${\rm PGFd}M\leq n$. Thus one has
${\rm Gpd}M\leq{\rm PGFd}M\leq{\rm pd}M;\ \ \ \ \ {\rm fd}M\leq{\rm pd}M;\ \ \ \ \ {\rm Gfd}M\leq{\rm min}\{{\rm PGFd}M,\ {\rm fd}M\}$
and
${\mathcal{F}_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{F}^{<\infty}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{GF}_{n}}$${\mathcal{P}^{<\infty}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{GF}^{<\infty}}$${\mathcal{PGF}_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{GP}_{n}}$${\mathcal{PGF}^{<\infty}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{GP}^{<\infty}.}$
1.2. Cotorsion pairs in exact categories
For exact category we refer to [26] and [21]. An exact category is weakly idempotent complete, if any splitting monomorphism has a cokernel; or equivalently, any splitting epimorphism has a kernel. For more equivalent descriptions see [7, 7.2, 7.6]. Any full subcategory of an abelian category which is closed under extensions and direct summands is a weakly idempotent complete exact category, with the natural exact structure, but it is not an abelian category, in general. The weakly idempotent complete exact categories considered in this paper are $\mathcal{PGF}_{n}$, $\mathcal{GP}_{n}$ and $\mathcal{GF}_{n}$, for each non-negative integer $n$, and $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$ and $\mathcal{GF}^{<\infty}$. They are not abelian in general.
For a class $\mathcal{C}$ of objects of exact category $\mathcal{A}$, let $\mathcal{C}^{\perp}$ be the class of objects $X$ with ${\rm Ext}^{1}_{\mathcal{A}}(C,X)=0,\ \forall\ C\in\mathcal{C}.$ Similarly for ${}^{\perp}\mathcal{C}$. A pair $(\mathcal{C},\mathcal{F})$ of classes of objects of $\mathcal{A}$ is a cotorsion pair, if $\mathcal{C}^{\perp}=\mathcal{F}$ and $\mathcal{C}=\ ^{\perp}\mathcal{F}$. For a cotorsion pair $(\mathcal{C},\mathcal{F})$,  $\mathcal{C}\cap\mathcal{F}$ is called its kernal.
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is complete, if for any object $X\in\mathcal{A}$, there are admissible exact sequences
$0\longrightarrow F\longrightarrow C\longrightarrow X\longrightarrow 0,\quad\text{and}\quad 0\longrightarrow X\longrightarrow F^{\prime}\longrightarrow C^{\prime}\longrightarrow 0,$
[…]
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$ for a set $\mathcal{S}$. Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([12, Theorem 10]; also [16, Theorem 6.11(b)]: but in [16] “cogenerated” is called “generated”). This result has been generalized to Grothendieck category with enough projective objects in [19, Theorem 2.4].
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is hereditary, if $\mathcal{C}$ is closed under the kernel of deflations, and $\mathcal{F}$ is closed under the cokernel of inflations. For a cotorsion pair $(\mathcal{C},\mathcal{F})$ in an abelian category $\mathcal{A}$ with enough projective objects and injective objects, these two conditions of the heredity are equivalent; and they are equivalent to ${{\rm Ext}}^{2}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$ and $F\in\mathcal{F};$ and also to ${{\rm Ext}}^{i}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$, $F\in\mathcal{F}$, and $i\geq 2$. See [13, Proposition 1.2.10]. This result holds also for a complete cotorsion pair in a weakly idempotent complete exact category. See [27, Lemma 6.17].
1.3. Model structures
A closed model structure, or simply, a model structure, on a category $\mathcal{M}$, is a triple  $(\operatorname{CoFib}$,  $\operatorname{Fib}$,  $\operatorname{Weq})$ of classes of morphisms, where the morphisms in the three classes are respectively called cofibrations, fibrations, and weak equivalences, satisfying Two out of three axiom, Retract axiom, Lifting axiom, and Factorization axiom. The morphisms in $\operatorname{CoFib}\cap\operatorname{Weq}$ (respectively, $\operatorname{Fib}\cap\operatorname{Weq}$) are called trivial cofibrations (respectively, trivial fibrations). For details we refer to [24], or [18]. In a model structure any two classes of $\operatorname{CoFib}$,   $\operatorname{Fib}$,   $\operatorname{Weq}$ uniquely determine the third one ([25]).
For a model structure $(\operatorname{CoFib}$,  $\operatorname{Fib}$,  $\operatorname{Weq})$ on category $\mathcal{M}$ with zero object, an object $X$ is trivial if $0\longrightarrow X$ is a weak equivalence, or, equivalently, $X\longrightarrow 0$ is a weak equivalence. It is cofibrant if $0\longrightarrow X$ is a cofibration, and it is fibrant if $X\longrightarrow 0$ is a fibration. An object is trivially cofibrant (respectively, trivially fibrant) if it is both trivial and cofibrant (respectively, fibrant). For a model structure on category $\mathcal{M}$ with zero object, Quillen’s homotopy category ${\rm Ho}(\mathcal{M})$ is the localization $\mathcal{M}[\operatorname{Weq}^{-1}]$. If $\mathcal{M}$ is an additive category, then ${\rm Ho}(\mathcal{M})$ is a pretriangulated category in the sense of [5]. It is not necessarily a triangulated category, in general.
A model structure on an exact category is exact ([14, 3.1]), if cofibrations are exactly inflations with cofibrant cokernel, and fibrations are exactly deflations with fibrant kernel. In this case, trivial cofibrations are exactly inflations with trivially cofibrant cokernel, and trivial fibrations are exactly deflations with trivially fibrant kernel. If $\mathcal{A}$ is an abelian category, then an exact model structure on $\mathcal{A}$ is exactly an abelian model structure in [19]. An exact model structure is projective if each object is fibrant, or equivalently, each trivially cofibrant object is projective ([14, 4.5, 4.6]). An exact model structure is trivial if each object is a trivial object. The homotopy category of a model structure is zero if and only if it is a trivial model structure ([18, Theorem 1.2.10(iv)]). We are only interested in exact model structures which are not trivial.
A Hovey triple ([19], [14]) in an exact category $\mathcal{A}$ is a triple  $(\mathcal{C},\mathcal{F},\mathcal{W})$  of classes of objects such that   $\mathcal{W}$ is thick in $\mathcal{A}$ (i.e., $\mathcal{W}$ is closed under direct summands, and if two out of three terms in an admissible exact sequence are in $\mathcal{W}$, then so is the third one); and that both $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})$ and  $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})$   are complete cotorsion pairs. We stress that in this paper a Hovey triple is written in the order  $(\mathcal{C},\mathcal{F},\mathcal{W})$, rather than  $(\mathcal{C},\mathcal{W},\mathcal{F})$.
M. Hovey [19] has established a one-one correspondence between abelian model structures and the Hovey triples in an abelian category. This has been extended to weakly idempotent complete exact category in J. Gillespie [14, Theorem 3.3] (see also [27, Theorem 6.9]). Namely, there is a one-to-one correspondence between exact model structures and the Hovey triples in weakly idempotent complete exact category $\mathcal{A}$, given by
$({\rm CoFib},\ {\rm Fib},\ {\rm Weq})\mapsto(\mathcal{C},\ \mathcal{F},\ \mathcal{W})$
where  $\mathcal{C}=\{\mbox{cofibrant objects}\},\ \ \mathcal{F}=\{\mbox{fibrant objects}\},\ \ \mathcal{W}=\{\mbox{trivial objects}\}$, with the inverse   $(\mathcal{C},\ \mathcal{F},\ \mathcal{W})\mapsto({\rm CoFib},\ {\rm Fib},\ {\rm Weq}),$ where
${\rm CoFib}=\{\mbox{inflation with cokernel in}\ \mathcal{C}\},\ \ \ {\rm Fib}=\{\mbox{deflation with kernel in}\ \mathcal{F}\},$
${\rm Weq}=\{pi\mid i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{C}\cap\mathcal{W},\ p\ \mbox{is a deflation,}\ {\rm Ker}p\in\mathcal{F}\cap\mathcal{W}\}.$
Thus, we identify a Hovey triple with an exact model structure, in weakly idempotent complete exact category. A Hovey triple $(\mathcal{C},\mathcal{F},\mathcal{W})$ is hereditary if $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})$ and  $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})$   are hereditary cotorsion pairs. An advantage of a hereditary Hovey triple is that the homotopy category of the corresponding exact model structure is just the stable category of a Frobenius category, and hence triangulated. See Lemma 1.3 below. Gillespie [15, Theorem 1.1] has constructed all the hereditary Hovey triples in an abelian category via two compatible complete cotorsion pairs.
1.4. Homotopy categories of exact model structures
The following observation is due to J. Št́ovíček, which shows that the class of weak equivalences of an exact model structure is uniquely determined only by the class of trivial objects, and hence its homotopy category is uniquely determined only by $\mathcal{W}$. This is highly non-trivial.
A morphism $f$ in exact category is admissible if $f=ip$, where $p$ is a deflation and $i$ is an inflation. In this case, ${\rm Ker}f$ and ${\rm Coker}f$ exist, and ${\rm Ker}f={\rm Ker}p$ and ${\rm Coker}f={\rm Coker}i$.
Lemma 1.1.
Let $\mathcal{A}$ be a weakly idempotent complete exact category, and $(\operatorname{CoFib},\ \operatorname{Fib},\ \operatorname{Weq})$ an exact model structure on $\mathcal{A}$ given by Hovey triple  $(\mathcal{C},\ \mathcal{F},\ \mathcal{W})$. Then
$(1)$   ([27, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
$(2)$  Put  $\mathcal{S}=\{f\in{\rm Mor}(\mathcal{A})\ |\ f\ \text{is admissible},\ {\rm Ker}f\in\mathcal{W},\ {\rm Coker}f\in\mathcal{W}\}.$ Then
$\operatorname{Weq}=\mathcal{S}\circ\mathcal{S}:=\{f\in{\rm Mor}(\mathcal{A})\ |\ f=f_{2}f_{1},\ f_{i}\in\mathcal{S},\ i=1,2\}.$
In particular, $\mathcal{S}\subseteq\operatorname{Weq}$.
Proof.
(2)  Since this assertion seems to be not available in the references, we include a proof. The inclusion $\operatorname{Weq}\subseteq\mathcal{S}\circ\mathcal{S}$ follow from (1); or alternatively, any weak equivalence $w$ has a decomposition $w=pi$ where $i$ is a trivial cofibration and $p$ is a trivial fibration. It remains to prove $\mathcal{S}\circ\mathcal{S}\subseteq\operatorname{Weq}$. It suffices to prove $\mathcal{S}\subseteq\operatorname{Weq}$. Let $f$ be an admissible morphism with ${\rm Ker}f\in\mathcal{W}$ and ${\rm Coker}f\in\mathcal{W}$. Then $f=pi$, where $i$ is a trivial cofibration and $p$ is a fibration. Thus $i$ is an inflation with ${\rm Coker}i\in\mathcal{C}\cap\mathcal{W}$ and $p$ is a deflation with ${\rm Ker}p\in\mathcal{F}$. By [7, Proposition 8.11], there is an admissible exact sequence
$0\longrightarrow{\rm Ker}f\longrightarrow{\rm Ker}p\longrightarrow{\rm Coker}i\longrightarrow{\rm Coker}f\longrightarrow 0$
Then one gets admissible exact sequences
$0\longrightarrow{\rm Ker}f\longrightarrow{\rm Ker}p\longrightarrow M\longrightarrow 0\ \ \ \ \ \text{and}\ \ \ 0\longrightarrow M\longrightarrow{\rm Coker}i\longrightarrow{\rm Coker}f\longrightarrow 0$
Since ${\rm Coker}i\in\mathcal{W}$ and ${\rm Coker}f\in\mathcal{W}$, $M\in\mathcal{W}$. Since $M\in\mathcal{W}$ and ${\rm Ker}f\in\mathcal{W}$, ${\rm Ker}p\in\mathcal{W}$. Thus $p$ is a deflation with ${\rm Ker}p\in\mathcal{F}\cap\mathcal{W}$, and hence $p$ is a trivial fibration. In particular, $p$ is a weak equivalence. Therefore $f=pi\in\operatorname{Weq}$. ∎
Corollary 1.2.
(Št́ovíček)  Let $\mathcal{A}$ be a weakly idempotent complete exact category, $(\mathcal{C},\mathcal{F},\mathcal{W})$ and $(\mathcal{C}^{\prime},\mathcal{F}^{\prime},\mathcal{W}^{\prime})$ Hovey triples in $\mathcal{A}$. If $W=W^{\prime}$, then the homotopy categories of the two exact model structures on $\mathcal{A}$ are the same.
Proof.
This follows from Lemma 1.1 and the definition of the homotopy category. ∎
For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([27, 6.15]; compare [18, 6.2], or, cofibration sequences and fibration sequences in [24]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5]. The following nice result can be explicitly founded in Št́ovíček [27, Theorem 6.21]. See also Gillespie [14, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
Theorem 1.3.
Let $\mathcal{A}$ be a weakly idempotent complete exact category, and $(\mathcal{C},\ \mathcal{F},\ \mathcal{W})$ a hereditary Hovey triple in $\mathcal{A}$. Then  $\mathcal{C}\cap\mathcal{F}$ is a Frobenius category, with the induced exact structure;  and $\mathcal{C}\cap\mathcal{F}\cap\mathcal{W}$ is the class of projective-injective objects of $\mathcal{C}\cap\mathcal{F}$.
Every cofiber sequence in ${\rm Ho}(\mathcal{A})$ is isomorphic in ${\rm Ho}(\mathcal{A})$ to a distinguished triangle in the stable category $(\mathcal{C}\cap\mathcal{F})/(\mathcal{C}\cap\mathcal{F}\cap\mathcal{W});$ and conversely every distinguished triangle in the stable category is isomorphic in ${\rm Ho}(\mathcal{A})$ to a cofiber sequence. Thus ${\rm Ho}(\mathcal{A})$ is a triangulated category and the composition $\mathcal{C}\cap\mathcal{F}\hookrightarrow\mathcal{A}\longrightarrow{\rm Ho}(\mathcal{A})$ induces a triangle equivalence
${\rm Ho}(\mathcal{A})\cong(\mathcal{C}\cap\mathcal{F})/(\mathcal{C}\cap\mathcal{F}\cap\mathcal{W}).$
Moreover, every conflation $0\longrightarrow X\stackrel{{\scriptstyle u}}{{\longrightarrow}}Y\stackrel{{\scriptstyle v}}{{\longrightarrow}}Z\longrightarrow 0$ in $\mathcal{A}$ yields a cofiber sequence $X\stackrel{{\scriptstyle u}}{{\longrightarrow}}Y\stackrel{{\scriptstyle v}}{{\longrightarrow}}Z\longrightarrow\Sigma X$ in ${\rm Ho}(\mathcal{A});$ and each cofiber sequence in ${\rm Ho}(\mathcal{A})$ is obtained in this way.
Corollary 1.4.
(Št́ovíček)  Let $\mathcal{A}$ be a weakly idempotent complete exact category. Let $(\mathcal{C},\mathcal{F},\mathcal{W})$ and $(\mathcal{C}^{\prime},\mathcal{F}^{\prime},\mathcal{W}^{\prime})$ be two hereditary Hovey triples in $\mathcal{A}$. If $\mathcal{W}=\mathcal{W}^{\prime}$, then
$(\mathcal{C}\cap\mathcal{F})/(\mathcal{C}\cap\mathcal{F}\cap\mathcal{W})\cong(\mathcal{C}^{\prime}\cap\mathcal{F}^{\prime})/(\mathcal{C}^{\prime}\cap\mathcal{F}^{\prime}\cap\mathcal{W}^{\prime})$
as triangulated categories.
Proof.
By Corollary 1.2, one only needs to show that the two triangulations in the same homotopy category also coincide. This is true, since by Theorem 1.3, the triangulation is determined by the cofiber sequences, which are by definition uniquely determined by conflations in $\mathcal{A}$ and by $\mathcal{W}$. See [27, Definition 6.15]. ∎
2. Abelian model structures arising from $\mathcal{PGF}_{n}$
The following remarkable result is due to Šaroch and Št́ovíček [28].
Lemma 2.1.
([28, Theorem 4.9])  Let $R$ be a ring. Then $\mathcal{PGF}^{\perp}$ is thick; and $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair with kernel $\mathcal{P}$, and cogenerated by a set.
The following interesting theorem can be found in E. E. Enochs and O. M. G. Jenda [11].
Lemma 2.2.
([11, Theorem 7.4.6])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete and hereditary cotorsion pair, and cogenerated by a set.
The PGF version of Auslander - Buchweitz Theorem ([2, Theorem 1.1]; see also [17, Theorem 2.10]) can be found in G. Dalezios and I. Emmanouil [8].
Lemma 2.3.
([8, Proposition 3.1])  Let $R$ be a ring, and $M$ an $R$-module with ${\rm PGFd}M=n$. Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{PGF}$ and ${\rm pd}K=n-1$. (normal-(((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{PGF}$-approximation of $M$.
The following fact can be found in G. Dalezios, I. Emmanouil [8].
Lemma 2.4.
([8, Proposition 2.4])  Let $R$ be a ring, and $0\longrightarrow M_{1}\longrightarrow M_{2}\longrightarrow M_{3}\longrightarrow 0$ an exact sequence. Then
${\rm(i)}$   ${\rm PGFd}M_{1}\leq{\rm max}\{{\rm PGFd}M_{2},{\rm PGFd}M_{3}\};$
${\rm(ii)}$   ${\rm PGFd}M_{2}\leq{\rm max}\{{\rm PGFd}M_{1},{\rm PGFd}M_{3}\};$
${\rm(iii)}$   ${\rm PGFd}M_{3}\leq{\rm max}\{{\rm PGFd}M_{1}+1,{\rm PGFd}M_{2}\}$.
The following result describes the modules of finite PFG dimension $\leq n$. It is a theorem of Šaroch - Št́ovíček type: [28, Theorem 4.11] is the similar result, for Gorenstein flat modules.
Theorem 2.5.
Let $R$ be a ring, $n$ a non-negative integer, and $M$ an $R$-module. Then $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n};$ and the following are equivalent:normal-:::
(1)  $M\in\mathcal{PGF}_{n}.$
(2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{R}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{PGF}$ and $K\in\mathcal{P}_{n-1}$. (normal-(((If $n=0$, then $K=0.)$
(3)  $\operatorname{Ext}_{R}^{1}(M,C)=0$ for every module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}.$
(4)  There is an exact sequence  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{PGF}$.
Proof.
$(1)\Longrightarrow(2)$:  Assume that $M\in\mathcal{PGF}_{n}.$ The case for $n=0$ is trivial. Assume that $n\geq 1$. By Lemma 2.3 there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{PGF}$ and $K\in\mathcal{P}_{n-1}$. It remains to see that this exact sequence remains exact after applying $\operatorname{Hom}_{R}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$. Since $G\in\mathcal{PGF}$, by definition there is an exact sequence $0\longrightarrow G\longrightarrow P\longrightarrow L\longrightarrow 0$ with $P$ projective and $L\in\mathcal{PGF}$. Take the pushout
[…]
It is clear that $H\in\mathcal{P}_{n}$. For any module $X\in\mathcal{P}_{n}^{\perp}$, in the following commutative diagram with exact rows
[…]
one has $\operatorname{Ext}_{R}^{1}(H,X)=0$ since $H\in\mathcal{P}_{n}$ and $X\in\mathcal{P}_{n}^{\perp}$. It follows that the map $\operatorname{Hom}_{R}(G,X)\longrightarrow\operatorname{Hom}_{R}(K,X)$ is surjective.
$(2)\Longrightarrow(3):$  Assume that $0\longrightarrow K\stackrel{{\scriptstyle f}}{{\longrightarrow}}G\longrightarrow M\longrightarrow 0$ is an exact sequence which remains exact after applying $\operatorname{Hom}_{R}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{PGF}$ and $K\in\mathcal{P}_{n-1}$. (If $n=0$ then $K=0$.) For any module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}$ one has ${\rm Ext}^{1}_{R}(G,C)=0$. Thus there is an exact sequence
${\rm Hom}_{R}(G,C)\stackrel{{\scriptstyle(f,C)}}{{\longrightarrow}}{\rm Hom}_{R}(K,C)\longrightarrow{\rm Ext}_{R}^{1}(M,C)\longrightarrow 0.$
But by the assumption the map $(f,C)$ is surjective, it follows that ${\rm Ext}_{R}^{1}(M,C)=0.$
$(3)\Longrightarrow(4)$:  Assume that $\operatorname{Ext}_{R}^{1}(M,C)=0$ for every module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}.$ By Lemma 2.1 one gets an exact sequence $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ with $L\in\mathcal{PGF}^{\perp}$ and $N\in\mathcal{PGF}$. Since $\operatorname{Ext}_{R}^{1}(N,C)=0$ for $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp},$ it follows that ${\rm Ext}^{1}_{R}(L,C)=0$ for $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}$.
By Lemma 2.2 there is an exact sequence $0\longrightarrow{\rm Ker}\pi\longrightarrow L^{\prime}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}L\longrightarrow 0$ with $L^{\prime}\in\mathcal{P}_{n}$ and ${\rm Ker}\pi\in\mathcal{P}_{n}^{\perp}$. Since $\mathcal{PGF}\subseteq\mathcal{GP}$, it follows that $L^{\prime}\in\mathcal{P}_{n}\subseteq\mathcal{GP}^{\perp}\subseteq\mathcal{PGF}^{\perp}$. Since both $L$ and $L^{\prime}$ are in $\mathcal{PGF}^{\perp}$, it follows from the thickness of $\mathcal{PGF}^{\perp}$ that ${\rm Ker}\pi\in\mathcal{PGF}^{\perp}$, and hence ${\rm Ker}\pi\in\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}$. Therefore ${\rm Ext}^{1}_{R}(L,{\rm Ker}\pi)=0$, and then $\pi$ splits and $L\in\mathcal{P}_{n}$.
$(4)\Longrightarrow(1)$: Assume that  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ is an exact sequence with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{PGF}$. Since $L\in\mathcal{P}_{n}\subseteq\mathcal{PGF}_{n}$ and $N\in\mathcal{PGF}\subseteq\mathcal{PGF}_{n}$, it follows $M\in\mathcal{PGF}_{n}$ (c.f. Lemma 2.4).
Finally, we show $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n}$. Since $\mathcal{P}_{n}\subseteq\mathcal{GP}^{\perp}\subseteq\mathcal{PGF}^{\perp}$ and $\mathcal{P}_{n}\subseteq\mathcal{PGF}_{n}$, thus $\mathcal{P}_{n}\subseteq\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}$. On the other hand, if $M\in\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}$, then by assertion (4) there is an exact sequence $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{PGF}$. Thus ${\rm Ext}^{1}_{R}(N,M)=0.$ Hence this exact sequence splits and $M\in\mathcal{P}_{n}$. ∎
When $n=0$, the following result coincides with [28, Theorem 4.9].
Corollary 2.6.
Let $R$ be a ring and $n$ a non-negative integer. Then
(1)  The pair  $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair in $R\mbox{-}{\rm Mod}$ with kernel $\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp}$. This cotorsion pair is cogenerated by a set.
(2)  The triple  $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple in $R\mbox{-}{\rm Mod};$ $\mathcal{PGF}_{n}\cap\mathcal{P}_{n}^{\perp}$ is a Frobenius category such that $\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp}$ is the class of projective-injective objects; and the homotopy category is the stable category $(\mathcal{PGF}_{n}\cap\mathcal{P}_{n}^{\perp})/(\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp})$, which is triangle equivalent to  $\mathcal{PGF}/\mathcal{P}.$
(3)  If $n\geq 1$, then the corresponding abelian model structure is projective if and only if $\mathcal{P}^{<\infty}=\mathcal{P}$.
Proof.
(1)  Put   $\mathcal{A}=\mathcal{PGF}_{n}$ and $\mathcal{B}=\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}$. By Theorem 2.5(3) one has $\mathcal{A}={}^{\perp}\mathcal{B}$ and $\mathcal{B}\subseteq\mathcal{A}^{\perp}$. On the other hand, one has
$\mathcal{P}_{n}\cup\mathcal{PGF}\subseteq\mathcal{PGF}_{n}\cup\mathcal{PGF}=\mathcal{PGF}_{n}.$
Thus
$\mathcal{A}^{\perp}=\mathcal{PGF}_{n}^{\perp}\subseteq(\mathcal{P}_{n}\cup\mathcal{PGF})^{\perp}=\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}=\mathcal{B}.$
Hence $\mathcal{B}=\mathcal{A}^{\perp}$ and $(\mathcal{A},\mathcal{B})$ is a cotorsion pair. It is hereditary, since $\mathcal{PGF}_{n}$ is closed under taking kernels of epimorphisms (cf. Lemma 2.4); or equivalently $\mathcal{P}_{n}^{\perp}$, and hence $\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}$, is closed under taking cokernels of monomorphisms.
To prove that it is complete, it suffices to show that it is cogenerated by a set. By [28, Theorem 4.9] (cf. Lemma 2.1), the cotorsion pair  $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is cogenerated by a set, say $T$. Also the cotorsion pair  $(\mathcal{P}_{n},\mathcal{P}_{n}^{\perp})$ is cogenerated by a set, say $B$. Then $\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}=B^{\perp}\cap T^{\perp}=(B\cup T)^{\perp}$, i.e., the cotorsion pair  $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is cogenerated by the set $B\cup T$.
By Theorem 2.5, $\mathcal{PGF}_{n}\cap\mathcal{PGF}^{\perp}=\mathcal{P}_{n}$, and hence the kernel of this cotorsion pair is $\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp}$.
(2)  Put $(\mathcal{C},\ \mathcal{F},\ \mathcal{W})=(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$. Then $\mathcal{W}=\mathcal{PGF}^{\perp}$ is thick (cf. Lemma 2.1). By Theorem 2.5 and Lemma 2.2, $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})=(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete and hereditary cotorsion pair; and $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})=(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair. By definition $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple in $R\mbox{-}{\rm Mod}$.
By Theorem 1.3, $\mathcal{C}\cap\mathcal{F}=\mathcal{PGF}_{n}\cap\mathcal{P}_{n}^{\perp}$ is a Frobenius category, and $\mathcal{C}\cap\mathcal{F}\cap\mathcal{W}=\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp}$ is its class of projective-injective objects; and the homotopy category is the stable category $(\mathcal{PGF}_{n}\cap\mathcal{P}_{n}^{\perp})/(\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp})$. Since $(\mathcal{PGF},\ R\mbox{-}{\rm Mod},\ \mathcal{PGF}^{\perp})$ is also a hereditary Hovey triple in $R\mbox{-}{\rm Mod}$, it follows from Corollary 1.4 that there is a triangle equivalence $(\mathcal{PGF}_{n}\cap\mathcal{P}_{n}^{\perp})/(\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
(3)  The class of fibrant objects of this model structure is  $\mathcal{P}_{n}^{\perp}$. By definition this model structure is projective if and only if $\mathcal{P}_{n}^{\perp}=R\mbox{-}{\rm Mod}.$ This is equivalent to say $\mathcal{P}_{n}=\ ^{\perp}(\mathcal{P}_{n}^{\perp})=\ ^{\perp}(R\mbox{-}{\rm Mod})=\mathcal{P}$, and this is equivalent to $\mathcal{P}^{<\infty}=\mathcal{P}$, if $n\geq 1$. ∎
Remark 2.7.
The completeness of the cotorsion pair in Corollary 2.6 can also be proved by the definition. See the proof of Theorem 3.4.
3. Abelian model structures arising from $\mathcal{GP}_{n}$
The following nice result is due to A. Beligiannis and I. Reiten [5].
Lemma 3.1.
([5, X, Theorem 2.4(iv)])  Let $A$ be an Artin algebra. Then  $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A$-Mod. Moreover, $\mathcal{GP}^{\perp}$ is thick in $A$-Mod (normal-(((this does not need the assumption of Artin algebra)normal-))).
Proof.
We include a proof for the thickness, for it is not mentioned in [5, X, Theorem 2.4(iv)]. Let $0\longrightarrow X\longrightarrow Y\longrightarrow Z\longrightarrow 0$ be an exact sequence, and $G\in\mathcal{GP}$. Then $G$ is a syzygy of a complete projective resolution $P^{\bullet}$. By exact sequence of complexes  $0\longrightarrow{\rm Hom}_{A}(P^{\bullet},X)\longrightarrow{\rm Hom}_{A}(P^{\bullet},Y)\longrightarrow{\rm Hom}_{A}(P^{\bullet},Z)\longrightarrow 0,$ one sees that any two of ${\rm Hom}_{A}(P^{\bullet},X),\ {\rm Hom}_{A}(P^{\bullet},Y)$,  ${\rm Hom}_{A}(P^{\bullet},Z)$ are exact, so is the third one. From this the thickness of $\mathcal{GP}^{\perp}$ follows. ∎
Auslander - Buchweitz Theorem is fundamental in Gorenstein homological algebra. See [2, Theorem 1.1], or more explicitly in [17, Theorem 2.10].
Lemma 3.2.
Let $R$ be a ring, and $M$ an $R$-module with ${\rm Gpd}M=n$. Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{GP}$ and ${\rm pd}K=n-1$. (normal-(((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{GP}$-approximation of $M$.
The following is another theorem of Šaroch - Št́ovíček type (compare Theorem 2.5 and [28, Theorem 4.11]).
Theorem 3.3.
Let $A$ be a ring, $n$ a non-negative integer, and $M$ an $A$-module. Then  $\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n};$ and one has the implications ${\rm(1)}\Longleftrightarrow{\rm(2)}\Longleftrightarrow{\rm(4)}\Longrightarrow{\rm(3)}$, where
(1)  $M\in\mathcal{GP}_{n}.$
(2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{A}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. (normal-(((If $n=0$, then $K=0.)$
(3)  $\operatorname{Ext}_{A}^{1}(M,C)=0$ for every module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}.$
(4)  There is an exact sequence  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{GP}$.
Moreover, if $A$ is an Artin algebra, then ${\rm(3)}\Longrightarrow{\rm(4)}$, and hence all the statements ${\rm(1)},\ {\rm(2)},\ {\rm(3)},\ {\rm(4)}$ are equivalent.
Proof.
It is clear that $\mathcal{P}_{n}\subseteq\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}$. Conversely, if $M\in\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}$, then by Lemma 3.2 there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. Since $K\in\mathcal{P}_{n-1}\subseteq\mathcal{GP}^{\perp}$ and $M\in\mathcal{GP}^{\perp}$, $G\in\mathcal{GP}^{\perp}$, and hence $G\in\mathcal{GP}\cap\mathcal{GP}^{\perp}=\mathcal{P}_{n}.$ Thus $M\in\mathcal{P}_{n}$.
$(1)\Longrightarrow(2)$:  Assume that $M\in\mathcal{GP}_{n}.$ The case for $n=0$ is trivial. Assume that $n\geq 1$. By Lemma 3.2 there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. We show that this sequence remains exact after applying $\operatorname{Hom}_{A}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$. Since $G$ is Gorenstein projective, by definition there is an exact sequence $0\longrightarrow G\longrightarrow P\longrightarrow G^{\prime}\longrightarrow 0$ with $P\in\mathcal{P}$ and $G^{\prime}\in\mathcal{GP}$. Take the pushout
[…]
Then $H\in\mathcal{P}_{n}$. For any module $X\in\mathcal{P}_{n}^{\perp}$, in the following commutative diagram with exact rows
[…]
one has $\operatorname{Ext}_{A}^{1}(H,X)=0$. It follows that the map $\operatorname{Hom}_{A}(G,X)\longrightarrow\operatorname{Hom}_{A}(K,X)$ is surjective.
$(2)\Longrightarrow(4)$:  The case when $n=0$ is trivial. If $n\geq 1$, by assumption there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ where $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. Also there is an exact sequence $0\longrightarrow G\longrightarrow P\longrightarrow N\longrightarrow 0$ where $P\in\mathcal{P}$ and $N\in\mathcal{GP}$. Take the pushout
[…]
Since $K\in\mathcal{P}_{n-1}$ and $P\in\mathcal{P}$, $L\in\mathcal{P}_{n}$. We are done.
$(4)\Longrightarrow(1)$: Assume that  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ is an exact sequence with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{GP}$. Since $L\in\mathcal{P}_{n}\subseteq\mathcal{GP}_{n}$ and $N\in\mathcal{GP}\subseteq\mathcal{GP}_{n}$, it follows from [17, Theorem 2.20] that $M\in\mathcal{GP}_{n}$.
$(2)\Longrightarrow(3):$  Assume that $0\longrightarrow K\stackrel{{\scriptstyle f}}{{\longrightarrow}}G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n}$ is an exact sequence which remains exact after applying $\operatorname{Hom}_{A}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$. For any module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$ one has ${\rm Ext}^{1}_{A}(G,C)=0$. Thus one has an exact sequence
${\rm Hom}_{A}(G,C)\stackrel{{\scriptstyle(f,C)}}{{\longrightarrow}}{\rm Hom}_{A}(K,C)\longrightarrow{\rm Ext}_{A}^{1}(M,C)\longrightarrow 0.$
But by the assumption the map $(f,C)$ is surjective, it follows that ${\rm Ext}_{A}^{1}(M,C)=0.$
$(3)\Longrightarrow(4)$:  For this, assume that $A$ is an Artin algebra. Assume that $\operatorname{Ext}_{A}^{1}(M,C)=0$ for every module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}.$ By Lemma 3.1 (This is the place we need the assumption that $A$ is an Artin algebra.), one gets an exact sequence $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ with $L\in\mathcal{GP}^{\perp}$ and $N\in\mathcal{GP}$. Since $\operatorname{Ext}_{A}^{1}(N,C)=0$ for $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp},$ it follows that ${\rm Ext}^{1}_{A}(L,C)=0$ for $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$.
By Lemma 2.2 one gets an exact sequence $0\longrightarrow{\rm Ker}\pi\longrightarrow L^{\prime}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}L\longrightarrow 0$ with $L^{\prime}\in\mathcal{P}_{n}$ and ${\rm Ker}\pi\in\mathcal{P}_{n}^{\perp}$. Since both $L$ and $L^{\prime}$ are in $\mathcal{GP}^{\perp}$, and since $\mathcal{GP}^{\perp}$ (cf. Lemma 3.1) is thick in $A$-Mod, it follows that ${\rm Ker}\pi\in\mathcal{GP}^{\perp}$, and hence ${\rm Ker}\pi\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$. Therefore ${\rm Ext}^{1}_{A}(L,{\rm Ker}\pi)=0$, and then $\pi$ splits and $L\in\mathcal{P}_{n}$. ∎
Since we do not know whether the cotorsion pair  $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is cogenerated by a set, the proof of the following theorem is in fact not the same as Corollary 2.6.
Theorem 3.4.
Let $A$ be an Artin algebra and $n$ a non-negative integer. Then
(1)  The pair  $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A\mbox{-}{\rm Mod}$ with kernel $\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp}$.
(2)  The triple  $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{GP}^{\perp})$ is a hereditary Hovey triple in $A\mbox{-}{\rm Mod};$ $\mathcal{GP}_{n}\cap\mathcal{P}_{n}^{\perp}$ is a Frobenius category such that $\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp}$ is its class of projective-injective objects; and the homotopy category is the stable category $(\mathcal{GP}_{n}\cap\mathcal{P}_{n}^{\perp})/(\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp})$, which is triangle equivalent to  $\mathcal{GP}/\mathcal{P}.$
Proof.
(1)  Put   $\mathcal{A}=\mathcal{GP}_{n}$ and $\mathcal{B}=\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$. By Theorem 3.3(3) one has $\mathcal{A}={}^{\perp}\mathcal{B}$ and $\mathcal{B}\subseteq\mathcal{A}^{\perp}$. On the other hand, one has  $\mathcal{P}_{n}\cup\mathcal{GP}\subseteq\mathcal{GP}_{n}\cup\mathcal{GP}=\mathcal{GP}_{n}.$ Thus
$\mathcal{A}^{\perp}=\mathcal{GP}_{n}^{\perp}\subseteq(\mathcal{P}_{n}\cup\mathcal{GP})^{\perp}=\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}=\mathcal{B}.$
Hence $\mathcal{B}=\mathcal{A}^{\perp}$ and $(\mathcal{A},\mathcal{B})$ is a cotorsion pair. It is hereditary, since $\mathcal{GP}_{n}$ is closed under taking kernels of epimorphisms (cf. [17, Theorem 2.20]); or equivalently $\mathcal{P}_{n}^{\perp}$, and hence $\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$, is closed under taking cokernels of monomorphisms.
To prove that it is complete, it suffices to show that for $M\in A\mbox{-}{\rm Mod}$, there is an exact sequence
$0\longrightarrow L\longrightarrow G\longrightarrow M\longrightarrow 0$
with $G\in\mathcal{GP}_{n}$ and $L\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$. Since $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete cotorsion pair in $A$-Mod (cf. Lemma 3.1), there is an exact sequence
$0\longrightarrow N\longrightarrow H\longrightarrow M\longrightarrow 0$
with $H\in\mathcal{GP}$ and $N\in\mathcal{GP}^{\perp}$. Since $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete cotorsion pair in $A\mbox{-}{\rm Mod}$ (cf. Lemma 2.2), there is an exact sequence
$0\longrightarrow N\longrightarrow L\longrightarrow C\longrightarrow 0$
with $L\in\mathcal{P}_{n}^{\perp}$ and $C\in\mathcal{P}_{n}$. Take the pushout
[…]
Since $N\in\mathcal{GP}^{\perp}$ and $C\in\mathcal{P}_{n}\subseteq\mathcal{GP}^{\perp}$, it follows that $L\in\mathcal{GP}^{\perp}$, and hence $L\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$. Since $H\in\mathcal{GP}\subseteq\mathcal{GP}_{n}$ and $C\in\mathcal{P}_{n}\subseteq\mathcal{GP}_{n}$, it follows that $G\in\mathcal{GP}_{n}$. We are done.
By Theorem 3.3,  $\mathcal{GP}_{n}\cap\mathcal{GP}^{\perp}=\mathcal{P}_{n},$ and hence the kernel of this cotorsion pair is $\mathcal{P}_{n}\cap\mathcal{P}_{n}^{\perp}$.
(2)  Put $(\mathcal{C},\ \mathcal{F},\ \mathcal{W})=(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{GP}^{\perp})$. Then $\mathcal{W}=\mathcal{GP}^{\perp}$ is thick, by Lemma 3.1. Since $\mathcal{GP}_{n}\cap\mathcal{GP}^{\perp}=\mathcal{P}_{n},$ it follows from Lemma 2.2 that $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})=(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete and hereditary cotorsion pair in $A\mbox{-}{\rm Mod}$. By (1), $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})=(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A\mbox{-}{\rm Mod}$. By definition $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{GP}^{\perp})$ is a hereditary Hovey triple in $A\mbox{-}{\rm Mod}$. The remaining assertion follows from Theorem 1.3 and Corollary 1.4, using hereditary Hovey triple $(\mathcal{GP},\ R\mbox{-}{\rm Mod},\ \mathcal{GP}^{\perp})$ in $R\mbox{-}{\rm Mod}$. ∎
Remark 3.5.
We do not need to consider the corresponding results for $\mathcal{GP}_{n}$ for Gorenstein ring $R$, by using the cortorsion pair  $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, since in this case one has $\mathcal{GP}=\mathcal{PGF}$ (cf. [20, Proposition 9]).
4. Several finitistic dimensions
4.1. Equalities on finitistic dimensions
For any non-negative integer $n$ and any ring $R$, we have seen in Corollary 2.6 that $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair in $R\mbox{-}{\rm Mod}$, and that $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple in $R\mbox{-}{\rm Mod}.$ Also, for Artin algebra $A$, we have seen in Theorem 3.4 that $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A\mbox{-}{\rm Mod}$, and that $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{GP}^{\perp})$ is a hereditary Hovey triple in $A\mbox{-}{\rm Mod}.$
For any ring $R$, Maaouy [22, Theorem A] proves that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R$-Mod.
It seems to be natural to ask whether $\mathcal{PGF}^{<\infty}$, $\mathcal{GF}^{<\infty}$ and $\mathcal{GF}^{<\infty}$ could be the left parts of cotorsion pairs in $R$-Mod. This problem reasonably leads to several finitistic dimensions.
The (left) finitistic projective dimension ${\rm Fpd}={\rm Fpd}(R)$ of a ring $R$ has been introduced by H. Bass ([3]), which is defined as
${\rm Fpd}={\rm sup}\{{\rm pd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm pd}M<\infty\}.$
Similarly, the (left) finitistic flat dimension ${\rm Ffd}={\rm Ffd}(R)$ of $R$ is defined as
${\rm Ffd}={\rm sup}\{{\rm fd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm fd}M<\infty\}.$
The (left) finitistic Gorenstein projective dimension ${\rm FGpd}={\rm FGpd}(R)$ ([17, 2.27]) is defined as
${\rm FGpd}={\rm sup}\{{\rm Gpd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm Gpd}M<\infty\}.$
The (left) finitistic Gorenstein flat dimension FGfd = FGfd$(R)$ ([17, 3.24]) is defined as
${\rm FGfd}={\rm sup}\{{\rm Gfd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm Gfd}M<\infty\}.$
Also, one can define the (left) finitistic PGF dimension of $R$ as
${\rm FPGFd}={\rm sup}\{{\rm PGFd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm PGFd}M<\infty\}.$
Using an argument of H. Holm, one can prove the following analogue of ([17, Theorem 2.28]).
Proposition 4.1.
Let $R$ be a ring. Then
(1)  ${\rm FPGFd}={\rm FGpd}={\rm Fpd}$.
(2)  If $R$ is right coherent and ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$.
(3)  If $R$ is right coherent, then ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
Proof.
(1)  [17, Theorem 2.28] claims that ${\rm FGpd}={\rm Fpd}$. Thus, it suffices to prove ${\rm FPGFd}={\rm Fpd}$.
Since $\mathcal{P}\subseteq\mathcal{PGF}\subseteq\mathcal{GP}$, it follows that ${\rm Gpd}M\leq{\rm PGFd}M\leq{\rm pd}M$ for any $R$-module.
If ${\rm pd}M<\infty$, then ${\rm Gpd}M={\rm pd}M$ (see [17, Proposition 2.27]), and hence ${\rm PGFd}M={\rm pd}M$ (see also [12, Corollary 3.7(i)]). It follows that ${\rm Fpd}\leq{\rm FPGFd}.$
If $0<{\rm PGFd}M=t<\infty$, then by [8, Proposition 3.1] (cf. Lemma 2.3) there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{PGF}$ and ${\rm pd}K=t-1.$ This implies that ${\rm FPGFd}\leq{\rm Fpd}+1$.
It remains to prove ${\rm FPGFd}\leq{\rm Fpd}.$ It suffices to show ${\rm FPGFd}\leq{\rm Fpd}$ under the assumption ${\rm FPGFpd}=m<\infty.$ (Otherwise ${\rm FPGFpd}=\infty,$ then by ${\rm FPGFd}\leq{\rm Fpd}+1$ one sees that ${\rm Fpd}=\infty$.) Take a module $M$ with ${\rm PGFd}M=m$, it suffices to find a module $L$ with ${\rm pd}L=m$. Thus one can assume that $m\geq 1$. By [8, Proposition 3.1] (cf. Lemma 2.3) there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{PGF}$ and ${\rm pd}K=m-1.$ Since $G\in\mathcal{PGF}$, $G$ is a submodule of a projective module $P$. Put $L=P/K$. Then ${\rm pd}L\leq m$. We claim that ${\rm pd}L=m$, and then we are done.
If $m=1$ then $K$ is projective. If $L$ is projective, then $0\longrightarrow K\longrightarrow P\longrightarrow L\longrightarrow 0$ splits, and hence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ splits. Thus $M\in\mathcal{PGF}$, i.e., ${\rm PGFd}M=0$. This contradicts the assumption ${\rm PGFd}M=m=1$. This proves ${\rm pd}L=1=m$.
If $m\geq 2$, then by ${\rm pd}K=m-1\geq 1$ there is a module $Z$ such that ${\rm Ext}^{m-1}_{R}(K,Z)\neq 0$. By the exact sequence ${\rm Ext}^{m-1}_{R}(P,Z)=0\longrightarrow{\rm Ext}^{m-1}_{R}(K,Z)\longrightarrow{\rm Ext}^{m}_{R}(L,Z)\longrightarrow 0={\rm Ext}^{m}_{R}(P,Z)$ one sees that ${\rm Ext}^{m}_{R}(L,Z)\neq 0$. Thus ${\rm pd}L=m$.
(2)   Assume that ${\rm fd}M=m<\infty$. Since $\mathcal{F}\subseteq\mathcal{GF}$, it follows that ${\rm Gfd}M\leq{\rm fd}M=m$. To prove ${\rm Gfd}M=m$, by [17, Theorem 3.14] it suffices to find an injective right $R$-module $I$ such that ${\rm Tor}^{R}_{m}(I,M)\neq 0$. Since ${\rm fd}M=m$, it follows that there is a right $R$-module $N$ such that ${\rm Tor}^{R}_{m}(N,M)\neq 0$. Consider an exact sequence $0\longrightarrow N\longrightarrow I\longrightarrow C\longrightarrow 0$ such that $I$ is an injective right $R$-module. Then by the exact sequence
${\rm Tor}^{R}_{m+1}(C,M)=0\longrightarrow{\rm Tor}^{R}_{m}(N,M)\longrightarrow{\rm Tor}^{R}_{m}(I,M)\longrightarrow{\rm Tor}^{R}_{m}(C,M)$
one sees that ${\rm Tor}^{R}_{m}(I,M)\neq 0$. We are done.
(3)  [17, Theorem 3.24] already claims that if $R$ is right coherent, then ${\rm FGfd}={\rm Ffd}$.
Since $\mathcal{F}\subseteq\mathcal{PGF}\subseteq\mathcal{GF}$, it follows that ${\rm Gfd}M\leq{\rm PGFd}M\leq{\rm fd}M$ for any $R$-module.
If ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$, by (2). It follows that ${\rm PGFd}M={\rm fd}M$. This implies that ${\rm Ffd}\leq{\rm FPGFd}.$ Together with (1) one gets ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
4.2. Abelian model structures arising from $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$, and $\mathcal{GF}^{<\infty}$
Corollary 4.2.
Let $R$ be a ring. Assume that the finitistic projective dimension ${\rm Fpd}$ of $R$ is finite, say ${\rm Fpd}=t$. Then
(1)  The pair  $(\mathcal{PGF}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})=(\mathcal{PGF}_{t},\ \mathcal{P}_{t}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{PGF}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})=(\mathcal{PGF}_{t},\ \mathcal{P}_{t}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
(2)  If $R$ is an Artin algebra, then  $(\mathcal{GP}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp}\cap\mathcal{GP}^{\perp})=(\mathcal{GP}_{t},\ \mathcal{P}_{t}^{\perp}\cap\mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{GP}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp},\ \mathcal{GP}^{\perp})=(\mathcal{GP}_{t},\ \mathcal{P}_{t}^{\perp},\ \mathcal{GP}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
(3)  If $R$ is right coherent, then  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
Proof.
(1)  By Proposition 4.1, one has ${\rm FGFd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}=t$, thus $\mathcal{PGF}^{<\infty}=\mathcal{PGF}_{t}$ and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  $\mathcal{GP}^{<\infty}=\mathcal{GP}_{t}$  and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  and $\mathcal{GF}^{<\infty}=\mathcal{GF}_{t}$  and $\mathcal{F}^{<\infty}=\mathcal{F}_{t}$. Then the assertions follow from Corollary 2.6, Theorem 3.4, and [22, Theorem A], respectively.∎
5. Induced complete cotorsion pairs in exact categories with applications
5.1. Induced complete cotorsion pairs in exact categories
The following observation will provide new complete cotorsion pairs in exact categories.
Theorem 5.1.
Let $\mathcal{A}$ be an abelian category, and $\mathcal{B}$ a full subcategory which is closed under extensions. Suppose that $(\mathcal{X},\ \mathcal{Y})$ is a complete cotorsion pair in $\mathcal{A}$ and $\mathcal{X}\subseteq\mathcal{B}$. If $\mathcal{B}$ is closed under the kernels of epimorphisms, then
$(1)$   $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in exact category $\mathcal{B}.$
$(2)$   If $(\mathcal{X},\ \mathcal{Y})$ is a hereditary cotorsion pair in $\mathcal{A}$, then $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a hereditary cotorsion pair in $\mathcal{B}$.
Proof.
$(1)$  It is clear that $\mathcal{B}$ is an exact category, and ${\rm Ext}^{1}_{\mathcal{B}}(X,Y)={\rm Ext}^{1}_{\mathcal{A}}(X,Y)=0$ for $X\in\mathcal{X}$ and $Y\in\mathcal{Y}\cap\mathcal{B}$. Thus $\mathcal{Y}\cap\mathcal{B}\subseteq\mathcal{X}^{\perp}\cap\mathcal{B}$ and $\mathcal{X}\subseteq\ ^{\perp}(\mathcal{Y}\cap\mathcal{B})\cap\mathcal{B}$. Let $M\in\mathcal{X}^{\perp}\cap\mathcal{B}$. Then it is clear that $M\in\mathcal{Y}\cap\mathcal{B}$. Thus $\mathcal{Y}\cap\mathcal{B}=\mathcal{X}^{\perp}\cap\mathcal{B}$. Let $M\in\ ^{\perp}(\mathcal{Y}\cap\mathcal{B})\cap\mathcal{B}$. Since $(\mathcal{X},\ \mathcal{Y})$ is a complete cotorsion pair in $\mathcal{A}$, there is an exact sequence $0\longrightarrow Y\longrightarrow X\longrightarrow M\longrightarrow 0$ with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$. Since $\mathcal{X}\subseteq\mathcal{B}$ and $\mathcal{B}$ is closed under the kernels of epimorphisms, $Y\in\mathcal{Y}\cap\mathcal{B}.$ Since $M\in\ ^{\perp}(\mathcal{Y}\cap\mathcal{B})\cap\mathcal{B}$,  ${\rm Ext}^{1}_{\mathcal{A}}(M,Y)=0$, thus the short exact sequence splits and $M\in\mathcal{X}.$
Thus $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a cotorsion pair in $\mathcal{B}.$ For $M\in\mathcal{B}$, as above there is an exact sequence $0\longrightarrow Y\longrightarrow X\longrightarrow M\longrightarrow 0$ with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}\cap\mathcal{B}$. By definition it is an admissible exact sequence in $\mathcal{B}.$ Also there is an exact sequence $0\longrightarrow M\longrightarrow Y^{\prime}\longrightarrow X^{\prime}\longrightarrow 0$ with $Y^{\prime}\in\mathcal{Y}$ and $X^{\prime}\in\mathcal{X}\subseteq\mathcal{B}$. Then $Y^{\prime}\in\mathcal{Y}\cap\mathcal{B}.$ Thus $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B}.$
$(2)$  Since $\mathcal{X}$ is closed under the kernels of epimorphisms in $\mathcal{A}$, $\mathcal{X}$ is closed under the kernels of deflations in $\mathcal{B}$. Since $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B},$ it follows that it is also hereditary (cf. [27, Lemma 6.17]; or Subsection 1.2). ∎
The dual of Theorem 5.1 is
Theorem 5.1’  Let $\mathcal{A}$ be an abelian category, and $\mathcal{B}$ an full subcategory which is closed under extensions. Suppose that $(\mathcal{X},\ \mathcal{Y})$ is a complete cotorsion pair in $\mathcal{A}$ and $\mathcal{Y}\subseteq\mathcal{B}$. If $\mathcal{B}$ is closed under the cokernels of monomorphisms, then
$(1)$   $(\mathcal{X}\cap\mathcal{B},\ \mathcal{Y})$ is a complete cotorsion pair in exact category $\mathcal{B}.$
$(2)$   If $(\mathcal{X},\ \mathcal{Y})$ is a hereditary cotorsion pair in $\mathcal{A}$, then $(\mathcal{X}\cap\mathcal{B},\ \mathcal{Y})$ is a hereditary cotorsion pair in $\mathcal{B}$.
5.2. Application 1: Exact model structures on $\mathcal{PGF}_{n}$
For each non-negative integer $n$, one has already known two chains of complete and hereditary cotorsion pairs $(\mathcal{X},\ \mathcal{Y})$ in $R$-Mod with $\mathcal{X}\subseteq\mathcal{PGF}_{n}$, namely $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}),\ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp})$, for each non-negative integer $m$ with $m\leq n$. See Lemma 2.2 and Corollary 2.6. Applying Theorem 5.1 one will get new chains of complete and hereditary cotorsion pairs, and exact model structures, on $\mathcal{PGF}_{n}$.
Note that $\mathcal{P}_{n}$ is not thick in $R$-Mod, in general.
Lemma 5.2.
Let $R$ be a ring and $n$ a non-negative integer. Then $\mathcal{P}_{n}$ is thick in $\mathcal{PGF}_{n}$ and in $\mathcal{GP}_{n}$.
Proof.
We only prove that $\mathcal{P}_{n}$ is thick in $\mathcal{PGF}_{n}$. Let $0\longrightarrow X\longrightarrow Y\longrightarrow Z\longrightarrow 0$ be an exact sequence in $\mathcal{PGF}_{n}$. It suffices to prove that if $X\in\mathcal{P}_{n}$ and $Y\in\mathcal{P}_{n}$ then $Z\in\mathcal{P}_{n}$. By $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n}$ (cf. Theorem 2.5) one has $Z\in\mathcal{P}_{n+1}\cap\mathcal{PGF}_{n}=\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n+1}\cap\mathcal{PGF}_{n}=\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n}.$
An alternative proof:  Since $Z\in\mathcal{PGF}_{n}$, by Lemma 2.3 there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow Z\longrightarrow 0$ with $G\in\mathcal{PGF}$ and $K\in\mathcal{P}_{n-1}$. (If $n=0$ then $K=0$.) The following proof holds also for $n=0$. Taking the pullback
[…]
Since $K\in\mathcal{P}_{n-1}$ and $Y\in\mathcal{P}_{n}$, it follows that $L\in\mathcal{P}_{n}$. Since $X\in\mathcal{P}_{n}$ and $L\in\mathcal{P}_{n}$, it follows that ${\rm pd}G<\infty.$ While $G\in\mathcal{PGF}$, $G$ is Gorenstein projective, and hence $G$ has to be projective (cf. [11, Proposition 10.2.3]). Thus $Z\in\mathcal{P}_{n}$. ∎
Corollary 5.3.
Let $R$ be a ring, $m$ and $n$ non-negative integers with $m\leq n$. Then
(1)  The pair  $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}_{n})$ is a complete and hereditary cotorsion pair in weakly idempotent complete exact category $\mathcal{PGF}_{n}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  The pair  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is a complete and hereditary cotorsion pair in $\mathcal{PGF}_{n}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(3)  The triple  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}_{n},\ \mathcal{P}_{n})$ is a hereditary Hovey triple in $\mathcal{PGF}_{n}$. In particular, $\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp}$ is a Frobenius category with $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$ as the class of projective-injective objects, and the corresponding homotopy category is  $(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
Proof.
Note that $\mathcal{PGF}_{n}$ is closed under extensions and the kernels of epimorphisms (cf. Lemma 2.4) and closed under direct summands (cf. [8, Proposition 2.3]). In particular, $\mathcal{PGF}_{n}$ is a weakly idempotent complete exact category.
(1)   By Lemma 2.2, $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, for each non-negative integer $m$. Since $m\leq n$, $\mathcal{P}_{m}\subseteq\mathcal{P}_{n}\subseteq\mathcal{PGF}_{n}.$ Then the assertion follows from Theorem 5.1.
(2)   By Corollary 2.6, $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, for each non-negative integer $m$. Since $m\leq n$, $\mathcal{PGF}_{m}\subseteq\mathcal{PGF}_{n}.$ Applying Theorem 5.1, one sees that $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n})$ is a complete and hereditary cotorsion pair in $\mathcal{PGF}_{n}$. By Theorem 2.5, $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n}$, thus this cotorsion pair is just  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}).$
Note that $\mathcal{PGF}_{m}\cap\mathcal{P}_{n}=\mathcal{P}_{m}.$ (This follows from $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n}$.) Thus the kernel of this cotorsion pair is $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(3)  Put $(\mathcal{C},\ \mathcal{F},\ \mathcal{W})=(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}_{n},\ \mathcal{P}_{n})$. By Lemma 5.2, $\mathcal{W}=\mathcal{P}_{n}$ is thick in $\mathcal{PGF}_{n}$.
Since $\mathcal{C}\cap\mathcal{W}=\mathcal{PGF}_{m}\cap\mathcal{P}_{n}=\mathcal{P}_{m}$, it follows from (1) that $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})=(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}_{n})$ is a complete and hereditary cotorsion pairs in $\mathcal{PGF}_{n}$. Also, $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})=(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is a complete and hereditary cotorsion pairs in $\mathcal{PGF}_{n}$, by (2). So the triple is a hereditary Hovey triple in $\mathcal{PGF}_{n}$. By Theorem 1.3, $\mathcal{C}\cap\mathcal{F}=\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp}$ is a Frobenius category, with
$\mathcal{C}\cap\mathcal{F}\cap\mathcal{W}=\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}=\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$
as the class of projective-injective objects, and the homotopy category is $(\mathcal{C}\cap\mathcal{F})/(\mathcal{C}\cap\mathcal{F}\cap\mathcal{W})=(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}),$ which is triangle equivalent to $\mathcal{PGF}/\mathcal{P}$, by Corollary 1.4. ∎
As we see, the exact model structure on $\mathcal{PGF}_{n}$ given in Corollary 5.3(3) is not trivial, and in general it is not projective.
5.3. Application 2: Exact model structures on $\mathcal{PGF}^{<\infty}$
It is clear that $\mathcal{P}_{m}\cap\mathcal{PGF}^{<\infty}=\mathcal{P}_{m}$. Also note that $\mathcal{PGF}_{m}\cap\mathcal{P}^{<\infty}=\mathcal{P}_{m}$. In fact, by $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{t}=\mathcal{P}_{t}$ for any non-negative integer (cf. Theorem 2.5) one has
$\mathcal{PGF}_{m}\cap\mathcal{P}^{<\infty}$ $=\bigcup\limits_{t\geq 0}(\mathcal{PGF}_{m}\cap\mathcal{P}_{t})=\bigcup\limits_{t\geq 0}(\mathcal{PGF}_{m}\cap\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{t})=\bigcup\limits_{t\geq 0}(\mathcal{P}_{m}\cap\mathcal{PGF}_{t})$
$=\mathcal{P}_{m}\cap\mathcal{PGF}^{<\infty}=\mathcal{P}_{m}.$
By the similar argument as in Corollary 5.3 one gets
Corollary 5.4.
Let $R$ be a ring and $m$ a non-negative integer. Then
(1)  The pair  $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{<\infty})$ is a complete and hereditary cotorsion pair in weakly idempotent complete exact category $\mathcal{PGF}^{<\infty}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  The pair  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}^{<\infty})$ is a complete and hereditary cotorsion pair in $\mathcal{PGF}^{<\infty}$ with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(3)  The triple  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{<\infty},\ \mathcal{P}^{<\infty})$ is a hereditary Hovey triple in $\mathcal{PGF}^{<\infty}$. The corresponding homotopy category is  $(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
When $m=0$, Corollary 5.4 has been obtained by Dalezios and Emmanouil [8, Theorem 4.1].
6. Exact model structures on $\mathcal{GP}_{n}$ and $\mathcal{GP}^{<\infty}$
6.1. Application 3: Exact model structures on $\mathcal{GP}_{n}$
For any non-negative integer $n$, applying Theorem 5.1 to $\mathcal{B}=\mathcal{GP}_{n}$ and the following two chains of complete and hereditary cotorsion pairs $(\mathcal{X},\ \mathcal{Y})$ in $R$-Mod with $\mathcal{X}\subseteq\mathcal{B}=\mathcal{GP}_{n}$:
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}),\ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}),$
where $0\leq m\leq n$ (see Lemma 2.2 and Corollary 2.6), one gets
Corollary 6.1.
Let $R$ be a ring, $m$ and $n$ non-negative integers with $m\leq n$. Then
(1)  The pair  $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n})$ is a complete and hereditary cotorsion pair in weakly idempotent complete exact category $\mathcal{GP}_{n}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  The pair  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ is a complete and hereditary cotorsion pair in $\mathcal{GP}_{n}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(3)  The triple  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ is a hereditary Hovey triple in $\mathcal{GP}_{n}$. The corresponding homotopy category is  $(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
Proof.
Note that $\mathcal{GP}_{n}$ is closed under extensions, the kernels of epimorphisms, and direct summands (cf. [17, Theorem 2.20, Proposition 2.19]). In particular, $\mathcal{GP}_{n}$ is a weakly idempotent complete exact category.
(1)  Applying Theorem 5.1 to $\mathcal{B}=\mathcal{GP}_{n}$ and the complete and hereditary cotorsion pair $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp})$ in $R$-Mod, one gets the assertion.
(2)  Applying Theorem 5.1 to $\mathcal{GP}_{n}$ and the complete and hereditary cotorsion pair $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp})$ in $R$-Mod, one gets the assertion. Since $\mathcal{PGF}_{m}\cap\mathcal{PGF}^{\perp}=\mathcal{P}_{m}$ (cf. Theorem 2.5), the kernel of this cotorsion pair is $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(3)  Since $\mathcal{PGF}^{\perp}$ is thick in $R$-Mod (cf. Lemma 2.1), $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}$ is thick in $\mathcal{GP}_{n}$. By (1) and (2),  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ is a hereditary Hovey triple in $\mathcal{GP}_{n}$. The last assertion follows from Theorem 1.3 and Corollary 1.4, by using that $(\mathcal{PGF},\ \mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ is also a hereditary Hovey triple in $\mathcal{GP}_{n}$. ∎
6.2. A new chain of cotorsion pairs and exact model structures on $\mathcal{GP}_{n}$
It is quite surprising that for $0\leq m\leq n$,  $(\mathcal{GP}_{m},\mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n})=(\mathcal{GP}_{m},\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is always a complete and hereditary cotorsion pair in $\mathcal{GP}_{n}$ (see Theorem 6.3), since in general we don’t know whether $(\mathcal{GP}_{m},\mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{\perp})$ is a cotorsion pair in $R\mbox{-}{\rm Mod}$ for an arbitrary ring $R$. Thus, Theorem 6.3 is not an application of Theorem 5.1 and Theorem 3.4. (Note that Theorem 3.4 needs the assumption of Artin algebra.)
Lemma 6.2.
Let $R$ be a ring. Then ${}^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{P}_{n}=\mathcal{P}_{m}$ in $R\mbox{-}{\rm Mod}$ for $0\leq m\leq n$.
Proof.
Clearly $\mathcal{P}_{m}\subseteq\ ^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{P}_{n}$. For the other hand, let $M\in\ ^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{P}_{n}$. Since  $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp})$ is a complete cotorsion pair (cf. Lemma 2.2), there is an exact sequence $0\longrightarrow N\longrightarrow H\longrightarrow M\longrightarrow 0$ with $H\in\mathcal{P}_{m}$ and $N\in\mathcal{P}_{m}^{\perp}$. Since $\mathcal{P}_{n}$ is closed under kernels of epimorphisms, $N\in\mathcal{P}_{n}$. Thus $N\in\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$ and this exact sequence splits. It follows that $M\in\mathcal{P}_{m}$. ∎
Theorem 6.3.
Let $R$ be a ring, $m$ and $n$ non-negative integers with $m\leq n$. Then
(1)  The pair $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is a complete and hereditary cotorsion pair in $\mathcal{GP}_{n}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  The triple  $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{P}_{n})$ is a hereditary Hovey triple in $\mathcal{GP}_{n}$. The corresponding homotopy category is  $(\mathcal{GP}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{GP}/\mathcal{P}.$
Proof.
(1)  To prove that $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is a cotorsion pair in $\mathcal{GP}_{n}$, by definition it suffices to prove that $\mathcal{GP}_{m}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$ and  ${}^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{GP}_{n}=\mathcal{GP}_{m}$.
Since $\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\subseteq\mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{\perp}$, it follows from Theorem 3.3(3) that
${\rm Ext}^{1}_{\mathcal{GP}_{n}}(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})={\rm Ext}^{1}_{R}(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})=0.$
It follows that $\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\subseteq\mathcal{GP}_{m}^{\perp}\cap\mathcal{GP}_{n}$ and $\mathcal{GP}_{m}\subseteq\ ^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{GP}_{n}$.
Since $\mathcal{GP}_{m}^{\perp}\subseteq\mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{\perp}$ and $\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$(cf. Theorem 3.3), it follows that $\mathcal{GP}_{m}^{\perp}\cap\mathcal{GP}_{n}\subseteq\mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$, and hence $\mathcal{GP}_{m}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$.
It remains to prove ${}^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{GP}_{n}\subseteq\mathcal{GP}_{m}$. Thus one can assume that $n\geq 1$. Let $M\in\ ^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{GP}_{n}$. By Lemma 3.2 there is an admissible exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. Since $G$ is Gorenstein projective, by definition there is an admissible exact sequence $0\longrightarrow G\longrightarrow P\longrightarrow G^{\prime}\longrightarrow 0$ with $P\in\mathcal{P}$ and $G^{\prime}\in\mathcal{GP}$. Take the pushout
[…]
Then $H\in\mathcal{P}_{n}$. Since $G^{\prime}\in\mathcal{GP}\subseteq\ ^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)$ and the class ${}^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)$ is closed under extensions, it follows that $H\in\ ^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)$. Thus $H\in\ ^{\perp}\left(\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}\right)\cap\mathcal{P}_{n}=\mathcal{P}_{m}$ (cf. Lemma 6.2), and hence $M\in\mathcal{GP}_{m}$ by Theorem 3.3 $(4)\Longrightarrow(1)$.
Up to now we have proved that $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is a cotorsion pair in $\mathcal{GP}_{n}$. Now we prove that it is complete. By Lemma 3.2, there is an admissible exact sequence
$0\longrightarrow N\longrightarrow H\longrightarrow M\longrightarrow 0$
with $H\in\mathcal{GP}$ and $N\in\mathcal{P}_{n-1}$. (If $n=0$ then $K=0$.) Since $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n})$ is a complete cotorsion pair in $\mathcal{GP}_{n}$, there is an admissible exact sequence
$0\longrightarrow N\longrightarrow L\longrightarrow C\longrightarrow 0$
with $L\in\mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n}$ and $C\in\mathcal{P}_{m}$. Take the pushout
[…]
Since $N\in\mathcal{P}_{n}$ and $C\in\mathcal{P}_{n}$, it follows that $L\in\mathcal{P}_{n}$, and hence $L\in\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$. Since $H\in\mathcal{GP}\subseteq\mathcal{GP}_{m}$ and $C\in\mathcal{P}_{m}\subseteq\mathcal{GP}_{m}$, it follows that $G\in\mathcal{GP}_{m}$. Thus we get an admissible exact sequence
$0\longrightarrow L\longrightarrow G\longrightarrow M\longrightarrow 0$
with $G\in\mathcal{GP}_{m}$ and $L\in\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$.
Now we show that for $M\in\mathcal{GP}_{n}$, there is an admissible exact sequence
$0\longrightarrow M\longrightarrow L\longrightarrow G\longrightarrow 0$
with $L\in\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$ and $G\in\mathcal{GP}_{m}$. By Theorem 3.3 $(1)\Longrightarrow(4)$, there is an admissible exact sequence
$0\longrightarrow M\longrightarrow N\longrightarrow H\longrightarrow 0$
with $N\in\mathcal{P}_{n}$ and $H\in\mathcal{GP}$. Since $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n})$ is a complete cotorsion pair in $\mathcal{GP}_{n}$ (cf. Corollary 6.1(1)), there is an admissible exact sequence
$0\longrightarrow N\longrightarrow L\longrightarrow C\longrightarrow 0$
with $L\in\mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n}$ and $C\in\mathcal{P}_{m}$. Take the pushout
[…]
Since $N\in\mathcal{P}_{n}$ and $C\in\mathcal{P}_{n}$, it follows that $L\in\mathcal{P}_{n}$, and hence $L\in\mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n}$. Since $H\in\mathcal{GP}\subseteq\mathcal{GP}_{m}$ and $C\in\mathcal{P}_{m}\subseteq\mathcal{GP}_{m}$, it follows that $G\in\mathcal{GP}_{m}$. We are done. That is, $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}_{n})$ is a complete cotorsion pair in $\mathcal{GP}_{n}$.
The heredity of this cotorsion pair follows from the fact that $\mathcal{GP}_{m}$ is closed under kernels of deflations in $\mathcal{GP}_{n}$. Since $\mathcal{GP}_{m}\cap\mathcal{P}_{n}=\mathcal{P}_{m},$ it follows that the kernel of this cotorsion pair is $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  Since $\mathcal{P}_{n}$ is thick in $\mathcal{GP}_{n}$ (cf. Lemma 5.2), it follows from (1) and Corollary 6.1(1) that  $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{P}_{n})$ is a hereditary Hovey triple in $\mathcal{GP}_{n}$. The last assertion follows from Theorem 1.3 and Corollary 1.4, by using that $(\mathcal{GP},\ \mathcal{GP}_{n},\ \mathcal{P}_{n})$ is also a hereditary Hovey triple in $\mathcal{GP}_{n}$. ∎
6.3. A sufficient and necessary condition for $\mathcal{PGF}=\mathcal{GP}$
Corollary 6.4.
Let $R$ be a ring. Then the following are equivalent:normal-:::
(1)  $\mathcal{PGF}=\mathcal{GP}.$
(2)  $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$ for any non-negative integer $n$.
(3)  $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$ for some non-negative integer $n$.
Proof.
$(1)\Longrightarrow(2)$:  By Theorem 3.3 one has $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$, for any non-negative integer $n$.
$(2)\Longrightarrow(3)$:  This is trivial.
$(3)\Longrightarrow(1)$:  Suppose that $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$ for some non-negative $n$. By Corollary 6.1(2) and Theorem 6.3(1), there are two cotorsion pairs $(\mathcal{PGF},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})=(\mathcal{PGF},\ \mathcal{P}_{n})$ and $(\mathcal{GP},\ \mathcal{P}_{n})$ in $\mathcal{GP}_{n}$. Thus $\mathcal{PGF}=\mathcal{GP}$. ∎
6.4. Exact model structures on $\mathcal{GP}^{<\infty}$
It is clear that $\mathcal{P}_{m}\cap\mathcal{GP}^{<\infty}=\mathcal{P}_{m}$. Also note that $\mathcal{GP}_{m}\cap\mathcal{P}^{<\infty}=\mathcal{P}_{m}$ (cf. [17, Proposition 2.27]). Similar to Corollary 6.1 and Theorem 6.3 one has
Corollary 6.5.
Let $R$ be a ring, and $m$ non-negative integers. Then
(1)  The pair  $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{<\infty})$ is a complete and hereditary cotorsion pair in weakly idempotent complete exact category $\mathcal{GP}^{<\infty}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  The pair  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GP}^{<\infty})$ is a complete and hereditary cotorsion pair in $\mathcal{GP}^{<\infty}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(3)  The pair  $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{P}^{<\infty})$ is a complete and hereditary cotorsion pair in $\mathcal{GP}^{<\infty}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(4)  The triple  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{<\infty},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}^{<\infty})$ is a hereditary Hovey triple in $\mathcal{GP}^{<\infty}$. The corresponding homotopy category is  $(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
(5)  The triple  $(\mathcal{GP}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}^{<\infty},\ \mathcal{P}^{<\infty})$ is a hereditary Hovey triple in $\mathcal{GP}^{<\infty}$. The corresponding homotopy category is  $(\mathcal{GP}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{GP}/\mathcal{P}.$
7. Exact model structures on $\mathcal{GF}_{n}$ and $\mathcal{GF}^{<\infty}$
7.1. Application 4: Exact model structures on $\mathcal{GF}_{n}$
The following result is due to R. EI Maaouy [22]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [28, Theorem 4.11].
Theorem 7.1.
([22, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent:normal-:::
(1)  $M\in\mathcal{GF}_{n}.$
(2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{R}(-,X)$ for $X\in\mathcal{F}_{n}^{\perp}$, where $G\in\mathcal{PGF}$, and $K$ is flat if $n=0$, and $K\in\mathcal{F}_{n-1}$ if $n\geq 1$.
(3)  $\operatorname{Ext}_{R}^{1}(M,C)=0$ for every module $C\in\mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}.$
(4)  There is an exact sequence  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ with $L\in\mathcal{F}_{n}$ and $N\in\mathcal{PGF}$.
The careful reader will find that the statement (2) is slightly stronger than [22, Theorem 3.4]: here $K\in\mathcal{F}_{n-1}$ if $n\geq 1$, rather than $K\in\mathcal{F}_{n}$ for all $n\geq 0$ as in [22, Theorem 3.4]; also, $X\in\mathcal{F}_{n}^{\perp}$ in (2), rather than $X\in\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp}$ as in [22, Theorem 3.4]. In this way, Theorem 7.1 coincides with the corresponding versions for modules in $\mathcal{PGF}_{n}$ as in Theorem 2.5, and for modules $\mathcal{GP}_{n}$ as in Theorem 3.3. We stress that the proof of [22, Theorem 3.4] completely holds for Theorem 7.1, thus we will not include a proof.
When $n=0$, the following result is [28, Theorem 4.9].
Theorem 7.2.
([22, Theorems A, B])  Let $R$ be a ring and $n$ a non-negative integer. Then
(1)  The pair  $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair in $R\mbox{-}{\rm Mod}$ with kernel $\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp}$. This cotorsion pair is cogenerated by a set.
(2)  The triple  $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple in $R\mbox{-}{\rm Mod};$ and the corresponding homotopy category is $(\mathcal{GF}_{n}\cap\mathcal{F}_{n}^{\perp})/(\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
Lemma 7.3.
([23, Theorem 3.4(2)])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, and cogenerated by a set.
The following result is a special case of D. Bennis, R. El Maaouy, J.R. García Rozas, L. Oyonarte [6], by taking $W=R$ in Proposition 7.9.
Lemma 7.4.
([6, Proposition 7.9])  Let $R$ be a ring, and $0\longrightarrow M_{1}\longrightarrow M_{2}\longrightarrow M_{3}\longrightarrow 0$ an exact sequence. Then
${\rm(i)}$   ${\rm Gfd}M_{1}\leq{\rm max}\{{\rm Gfd}M_{2},{\rm Gfd}M_{3}-1\};$
${\rm(ii)}$   ${\rm Gfd}M_{2}\leq{\rm max}\{{\rm Gfd}M_{1},{\rm Gfd}M_{3}\};$
${\rm(iii)}$   ${\rm Gfd}M_{3}\leq{\rm max}\{{\rm Gfd}M_{1}+1,{\rm Gfd}M_{2}\}$.
We specially appreciate the best upper bound given in (i), as for projective dimension, rather than a correct upper bound: ${\rm Gfd}M_{1}\leq{\rm max}\{{\rm Gfd}M_{2},{\rm Gfd}M_{3}\}.$ This is really needed in the proof of Theorem 7.1(2).
Lemma 7.5.
Let $R$ be a ring, and $M$ an $R$-module with ${\rm fd}M<\infty$.
(1)  ([9, Remark 1.5])  If $M\in\mathcal{GF}$, then $M$ is flat.
(2)  ([8])  If $M\in\mathcal{PGF}$, then $M$ is projective.
Proof.
The assertion (2) is implicitly contained in [8]. For convenience we include a justification. By [8, Proposition 2.5(ii)], ${\rm pd}M={\rm fd}M<\infty$. Since $M\in\mathcal{PGF}$, it follows that $M$ is Gorenstein projective ([28, Theorem 4.4]), with finite projective dimension, and hence $M$ is projective (cf. [11, Proposition 10.2.3]). An alternative proof: by Theorem 7.1 one has $M\in\mathcal{PGF}\cap\mathcal{F}_{n}=\mathcal{PGF}\cap\mathcal{GF}_{n}\cap\mathcal{PGF}^{\perp}=(\mathcal{PGF}\cap\mathcal{PGF}^{\perp})\cap\mathcal{GF}_{n}=\mathcal{P}\cap\mathcal{GF}_{n}=\mathcal{P}.$
Lemma 7.6.
Let $R$ be a ring and $n$ a non-negative integer. Then $\mathcal{F}_{n}$ is thick in $\mathcal{GF}_{n}$.
Proof.
If $n=0$ this in known in [8, Theorem 4.3], by using Lemma 7.5(1). Assume that $n\geq 1$. The proof is the similar as Lemma 5.2. Let $0\longrightarrow X\longrightarrow Y\longrightarrow Z\longrightarrow 0$ be an exact sequence in $\mathcal{GF}_{n}$. It suffices to prove that if $X\in\mathcal{F}_{n}$ and $Y\in\mathcal{F}_{n}$ then $Z\in\mathcal{F}_{n}$. By $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n}$ (cf. Theorem 7.1) one has $Z\in\mathcal{F}_{n+1}\cap\mathcal{GF}_{n}=\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n+1}\cap\mathcal{GF}_{n}=\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n}.$
For any non-negative integer $n$, we have four chains of complete and hereditary cotorsion pairs $(\mathcal{X},\ \mathcal{Y})$ in $R$-Mod with $\mathcal{X}\subseteq\mathcal{B}=\mathcal{GF}_{n}$:
$(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}),\ \ \ \ (\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}),\ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}),\ \ \ \ (\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{PGF}^{\perp})$
where $0\leq m\leq n$ (see Lemma 2.2, Lemma 7.3, Corollary 2.6, and Theorem 7.2). Applying Theorem 5.1 to $\mathcal{B}=\mathcal{GF}_{n}$, one gets the assertions (1) - (4) in the following corollary.
Corollary 7.7.
Let $R$ be a ring, $m$ and $n$ non-negative integers with $m\leq n$. Then
(1)  The pair  $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}_{n})$ is a complete and hereditary cotorsion pair in weakly idempotent complete exact category $\mathcal{GF}_{n}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  The pair  $(\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}_{n})$ is a complete and hereditary cotorsion pair in $\mathcal{GF}_{n}$, with kernel $\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp}$.
(3)  The pair  $(\mathcal{PGF}_{m},\ \mathcal{F}_{n}\cap\mathcal{P}_{m}^{\perp})$ is a complete and hereditary cotorsion pair in $\mathcal{GF}_{n}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(4)  The pair  $(\mathcal{GF}_{m},\ \mathcal{F}_{n}\cap\mathcal{F}_{m}^{\perp})$ is a complete and hereditary cotorsion pair in $\mathcal{GF}_{n}$, with kernel $\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp}$.
(5)  The triple  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}_{n},\ \mathcal{F}_{n})$ is a hereditary Hovey triple in $\mathcal{GF}_{n}$. The corresponding homotopy category is  $(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
(6)  The triple  $(\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}_{n},\ \mathcal{F}_{n})$ is a hereditary Hovey triple in $\mathcal{GF}_{n}$. The corresponding homotopy category is  $(\mathcal{GF}_{m}\cap\mathcal{F}_{m}^{\perp})/(\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
Proof.
Note that $\mathcal{GF}_{n}$ is closed under extensions (cf. Lemma 7.4(ii)), and closed under direct summands (cf. [6, Corollary 7.8(2)]). Thus $\mathcal{GF}_{n}$ is a weakly idempotent complete exact category. Also, $\mathcal{GF}_{n}$ is closed under the kernels of epimorphisms (cf. Lemma 7.4(i)): this is a requirement for the applications of Theorem 5.1.
(1)  Applying Theorem 5.1 to $\mathcal{B}=\mathcal{GF}_{n}$ and $(\mathcal{X},\ \mathcal{Y})=(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp})$, one gets (1).
(2)  Applying Theorem 5.1 to $\mathcal{GF}_{n}$ and $(\mathcal{X},\ \mathcal{Y})=(\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp})$, one gets (2).
(3)  Applying Theorem 5.1 to $\mathcal{GF}_{n}$ and $(\mathcal{X},\ \mathcal{Y})=(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp})$, one gets the complete and hereditary cotorsion pair $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n})$ in $\mathcal{GF}_{n}$. Since $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n}$ (cf. Theorem 7.1), this cotorsion pair is just $(\mathcal{PGF}_{m},\ \mathcal{F}_{n}\cap\mathcal{P}_{m}^{\perp})$.
Since $\mathcal{PGF}_{m}\cap\mathcal{F}_{n}=\mathcal{PGF}_{m}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{PGF}_{m}\cap\mathcal{PGF}^{\perp}=\mathcal{P}_{m}$ (cf. Theorem 2.5), the kernel of this cotorsion pair is $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(4)  Applying Theorem 5.1 to $\mathcal{GF}_{n}$ and $(\mathcal{X},\ \mathcal{Y})=(\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{PGF}^{\perp})$, one gets the complete and hereditary cotorsion pair $(\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n})$ in $\mathcal{GF}_{n}$. By $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n}$ (cf. Theorem 7.1), this cotorsion pair is just $(\mathcal{GF}_{m},\ \mathcal{F}_{n}\cap\mathcal{F}^{\perp}_{m})$.
Since $\mathcal{GF}_{m}\cap\mathcal{F}_{n}=\mathcal{GF}_{m}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{GF}_{m}\cap\mathcal{PGF}^{\perp}=\mathcal{F}_{m}$, the kernel of this cotorsion pair is $\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp}$.
(5)  By Lemma 7.6, $\mathcal{F}_{n}$ is thick in $\mathcal{GF}_{n}$.
Since $\mathcal{PGF}_{m}\cap\mathcal{F}_{n}=\mathcal{P}_{m}$, it follows from (1) and (3) that $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}_{n},\ \mathcal{F}_{n})$ is a hereditary Hovey triple in $\mathcal{GF}_{n}$. Thus, by Theorem 1.3, $\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp}$ is a Frobenius category, $\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp}\cap\mathcal{GF}_{n}\cap\mathcal{F}_{n}=\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$ is the class of its projective-injective objects, and the homotopy category of the corresponding exact model structure on $\mathcal{GF}_{n}$ is  $(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}).$ It is triangle equivalent to $\mathcal{PGF}/\mathcal{P},$ by using Corollary 1.4 and hereditary Hovey triple $(\mathcal{PGF},\ \mathcal{GF}_{n},\ \mathcal{F}_{n})$ in $\mathcal{GF}_{n}$.
(6)  By Lemma 7.6, $\mathcal{F}_{n}$ is thick in $\mathcal{GF}_{n}$.
Since $\mathcal{GF}_{m}\cap\mathcal{F}_{n}=\mathcal{F}_{m}$, it follows from (2) and (4) that  $(\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}_{n},\ \mathcal{F}_{n})$ is a hereditary Hovey triple in $\mathcal{GF}_{n}$. Thus, by Theorem 1.3, $\mathcal{GF}_{m}\cap\mathcal{F}_{m}^{\perp}$ is a Frobenius category, $\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp}$ is the class of its projective-injective objects, and the homotopy category of the corresponding exact model structure on $\mathcal{GF}_{n}$ is  $(\mathcal{GF}_{m}\cap\mathcal{F}_{m}^{\perp})/(\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp}).$ By (5), $(\mathcal{PGF},\ \mathcal{GF}_{n},\ \mathcal{F}_{n})$ is also a hereditary Hovey triple in $\mathcal{GF}_{n}$, with homotopy category $\mathcal{PGF}/\mathcal{P}.$ By Corollary 1.4, there is a triangle equivalence $(\mathcal{GF}_{m}\cap\mathcal{F}_{m}^{\perp})/(\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}$. ∎
When $n=m=0$, Corollary 7.7(3) and (5) has been obtained by G. Dalezios and I. Emmanouil [8, Theorem 4.3].
7.2. Application 5: Exact model structures on $\mathcal{GF}^{<\infty}$
Fact 7.8.
Let $R$ be a ring and $m$ a non-negative integer. Then
(1)  $\mathcal{PGF}^{\perp}\cap\mathcal{GF}^{<\infty}=\mathcal{F}^{<\infty}.$
(2)  $\mathcal{PGF}_{m}\cap\mathcal{F}^{<\infty}=\mathcal{P}_{m}.$
(3)   $\mathcal{GF}_{m}\cap\mathcal{F}^{<\infty}=\mathcal{F}_{m}.$
Proof.
One has $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{t}=\mathcal{P}_{t}$ (cf. Theorem 2.5) and $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{t}=\mathcal{F}_{t}$ (cf. Theorem 7.1), for any non-negative integer $t$.
(1)  One has $\mathcal{PGF}^{\perp}\cap\mathcal{GF}^{<\infty}=\bigcup\limits_{t\geq 0}(\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{t})=\bigcup\limits_{t\geq 0}\mathcal{F}_{t}=\mathcal{F}^{<\infty}.$
(2)  One has $\mathcal{PGF}_{m}\cap\mathcal{F}^{<\infty}=\bigcup\limits_{t\geq 0}(\mathcal{PGF}_{m}\cap\mathcal{F}_{t})=\bigcup\limits_{t\geq 0}(\mathcal{PGF}_{m}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{t})=\bigcup\limits_{t\geq 0}(\mathcal{P}_{m}\cap\mathcal{GF}_{t})=\mathcal{P}_{m}\cap\mathcal{GF}^{<\infty}=\mathcal{P}_{m}.$
(3)  One has $\mathcal{GF}_{m}\cap\mathcal{F}^{<\infty}=\bigcup\limits_{t\geq 0}(\mathcal{GF}_{m}\cap\mathcal{F}_{t})=\bigcup\limits_{t\geq 0}(\mathcal{GF}_{m}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{t})=\bigcup\limits_{t\geq 0}(\mathcal{F}_{m}\cap\mathcal{GF}_{t})=\mathcal{F}_{m}\cap\mathcal{GF}^{<\infty}=\mathcal{F}_{m}.$
By the similar argument as in Corollary 7.7 one has
Corollary 7.9.
Let $R$ be a ring and $m$ a non-negative integers. Then
(1)  The pair  $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}^{<\infty})$ is a complete and hereditary cotorsion pair in weakly idempotent complete exact category $\mathcal{GF}^{<\infty}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(2)  The pair  $(\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}^{<\infty})$ is a complete and hereditary cotorsion pair in $\mathcal{GF}^{<\infty}$, with kernel $\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp}$.
(3)  The pair  $(\mathcal{PGF}_{m},\ \mathcal{F}^{<\infty}\cap\mathcal{P}_{m}^{\perp})$ is a complete and hereditary cotorsion pair in $\mathcal{GF}^{<\infty}$, with kernel $\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp}$.
(4)  The pair  $(\mathcal{GF}_{m},\ \mathcal{F}^{<\infty}\cap\mathcal{F}_{m}^{\perp})$ is a complete and hereditary cotorsion pair in $\mathcal{GF}^{<\infty}$, with kernel $\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp}$.
(5)  The triple  $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GF}^{<\infty},\ \mathcal{F}^{<\infty})$ is a hereditary Hovey triple in $\mathcal{GF}^{<\infty}$. The corresponding homotopy category is  $(\mathcal{PGF}_{m}\cap\mathcal{P}_{m}^{\perp})/(\mathcal{P}_{m}\cap\mathcal{P}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
(6)  The triple  $(\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{GF}^{<\infty},\ \mathcal{F}^{<\infty})$ is a hereditary Hovey triple in $\mathcal{GF}^{<\infty}$. The corresponding homotopy category is  $(\mathcal{GF}_{m}\cap\mathcal{F}_{m}^{\perp})/(\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$

### v3

@@
 2020 Mathematics Subject Classification. Primary 16E30, 18N40; Secondary 16E10, 16E65, 16G50
-$*{}^{*}$ Corresponding author
+Corresponding author
 […]
@@
 Jenda [10], have got deep interests and wide applications in mathematics.
-A recent milestone is the introduction of projectively coresolved Gorenstein flat (PGF, for short) modules over an arbitrary ring $R$, by J.
+A recent important progress is the introduction of projectively coresolved Gorenstein flat (PGF, for short) modules over an arbitrary ring $R$, by J.
 Šaroch and J.
@@
 Beligiannis and I.
-Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A$-Mod, if $A$ is an Artin algebra.
+Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra.
 Šaroch and Št́ovíček [28, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$.
@@
 R.
-EI Maaouy [22, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
+El Maaouy [22, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
 $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
@@
 (We remind that, in this paper, a Hovey triple is written in the order  $(\mathcal{C},\mathcal{F},\mathcal{W})$, rather than  $(\mathcal{C},\mathcal{W},\mathcal{F})$, where $\mathcal{W}$ is the class of trivial objects.) Since all the hereditary abelian model structures in each chain have the same class of trivial objects, the homotopy categories in the same chain are triangle equivalent to each other.
-Thus the homotopy categories are the stable categories $\mathcal{PGF}/\mathcal{P}$ and $\mathcal{GP}/\mathcal{P}$.
+Thus the homotopy categories are the stable categories $\mathcal{PGF}/\mathcal{P}$ and $\mathcal{GP}/\mathcal{P}$, respectively.
 One may further ask whether $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$, and $\mathcal{GF}^{<\infty}$, can be the left parts of complete and hereditary cotorsion pairs in $R$-Mod, where $\mathcal{PGF}^{<\infty}$ (respectively, $\mathcal{GP}^{<\infty},\ \mathcal{GF}^{<\infty}$) is the class of $R$-modules of finite PGF (respectively, the Gorenstein projective, Gorenstein flat) dimension.
@@
 $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n}),\ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n});$
-and one chain of hereditary Hovey triples $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ in $\mathcal{GF}_{n}$.
+and one chain of hereditary Hovey triples $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ in $\mathcal{GP}_{n}$.
 See Corollary 6.1.
 In this way one gets double chains of complete and hereditary cotorsion pairs and hereditary Hovey triples, parameterized by $m$ and $n$.
-When $m=0=n$, this is G.
-Dalezios and I.
-Emmanouil [8, Theorem 4.3].
 All these exact model structures are not trivial, and not projective in general, with the same homotopy category $\mathcal{PGF}/\mathcal{P}$.
@@
 […]
-A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$ for a set $\mathcal{S}$.
+A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$, where $\mathcal{S}$ is a set of modules (or, equivalently, cogenerated by a single module).
 Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([12, Theorem 10]; also [16, Theorem 6.11(b)]: but in [16] “cogenerated” is called “generated”).
@@
 This result holds also for a complete cotorsion pair in a weakly idempotent complete exact category.
-See [27, Lemma 6.17].
+See [29, Lemma 6.17] (also [27]).
 1.3. Model structures
 A closed model structure, or simply, a model structure, on a category $\mathcal{M}$, is a triple  $(\operatorname{CoFib}$,  $\operatorname{Fib}$,  $\operatorname{Weq})$ of classes of morphisms, where the morphisms in the three classes are respectively called cofibrations, fibrations, and weak equivalences, satisfying Two out of three axiom, Retract axiom, Lifting axiom, and Factorization axiom.
-The morphisms in $\operatorname{CoFib}\cap\operatorname{Weq}$ (respectively, $\operatorname{Fib}\cap\operatorname{Weq}$) are called trivial cofibrations (respectively, trivial fibrations).
+The morphisms in ${\rm TCoFib}:=\operatorname{CoFib}\cap\operatorname{Weq}$ (respectively, ${\rm TFib}:=\operatorname{Fib}\cap\operatorname{Weq}$) are called trivial cofibrations (respectively, trivial fibrations).
 For details we refer to [24], or [18].
@@
 This has been extended to weakly idempotent complete exact category in J.
-Gillespie [14, Theorem 3.3] (see also [27, Theorem 6.9]).
+Gillespie [14, Theorem 3.3] (see also [29, Theorem 6.9]).
 Namely, there is a one-to-one correspondence between exact model structures and the Hovey triples in weakly idempotent complete exact category $\mathcal{A}$, given by
@@
 ${\rm CoFib}=\{\mbox{inflation with cokernel in}\ \mathcal{C}\},\ \ \ {\rm Fib}=\{\mbox{deflation with kernel in}\ \mathcal{F}\},$
-${\rm Weq}=\{pi\mid i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{C}\cap\mathcal{W},\ p\ \mbox{is a deflation,}\ {\rm Ker}p\in\mathcal{F}\cap\mathcal{W}\}.$
+${\rm Weq}=\{pi\mid i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{C}\cap\mathcal{W},\ p\ \mbox{is a deflation,}\ {\rm Ker}p\in\mathcal{F}\cap\mathcal{W}\}$
+and
+${\rm TCoFib}=\{\mbox{inflation with cokernel in}\ \mathcal{C}\cap\mathcal{W}\},\ \ \ \ {\rm TFib}=\{\mbox{deflation with kernel in}\ \mathcal{F}\cap\mathcal{W}\}.$
 Thus, we identify a Hovey triple with an exact model structure, in weakly idempotent complete exact category.
@@
 In this case, ${\rm Ker}f$ and ${\rm Coker}f$ exist, and ${\rm Ker}f={\rm Ker}p$ and ${\rm Coker}f={\rm Coker}i$.
+The following description of $\operatorname{Weq}$ in $(2)$ is similar to the multiplicative system determined by a Serre subcategory of an abelian category.
 Lemma 1.1.
@@
 Then
-$(1)$   ([27, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
+$(1)$   ([29, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
 $(2)$  Put  $\mathcal{S}=\{f\in{\rm Mor}(\mathcal{A})\ |\ f\ \text{is admissible},\ {\rm Ker}f\in\mathcal{W},\ {\rm Coker}f\in\mathcal{W}\}.$ Then
@@
 This follows from Lemma 1.1 and the definition of the homotopy category. ∎
-For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([27, 6.15]; compare [18, 6.2], or, cofibration sequences and fibration sequences in [24]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5].
-The following nice result can be explicitly founded in Št́ovíček [27, Theorem 6.21].
+For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([29, 6.15]; compare [18, 6.2], or, cofibration sequences and fibration sequences in [24]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5].
+The following nice result can be explicitly founded in Št́ovíček [29, Theorem 6.21].
 See also Gillespie [14, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
@@
 This is true, since by Theorem 1.3, the triangulation is determined by the cofiber sequences, which are by definition uniquely determined by conflations in $\mathcal{A}$ and by $\mathcal{W}$.
-See [27, Definition 6.15].
+See [29, Definition 6.15].
 ∎
@@
 Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{PGF}$ and ${\rm pd}K=n-1$.
-(normal-(((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{PGF}$-approximation of $M$.
+((((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{PGF}$-approximation of $M$.
 The following fact can be found in G. Dalezios, I. Emmanouil [8].
@@
 Theorem 2.5.
-Let $R$ be a ring, $n$ a non-negative integer, and $M$ an $R$-module. Then $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n};$ and the following are equivalent:normal-:::
+Let $R$ be a ring, $n$ a non-negative integer, and $M$ an $R$-module. Then $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n};$ and the following are equivalent::::
 (1)  $M\in\mathcal{PGF}_{n}.$
 (2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{R}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{PGF}$ and $K\in\mathcal{P}_{n-1}$.
-(normal-(((If $n=0$, then $K=0.)$
+((((If $n=0$, then $K=0.)$
 (3)  $\operatorname{Ext}_{R}^{1}(M,C)=0$ for every module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}.$
@@
 Then  $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A$-Mod.
-Moreover, $\mathcal{GP}^{\perp}$ is thick in $A$-Mod (normal-(((this does not need the assumption of Artin algebra)normal-))).
+Moreover, $\mathcal{GP}^{\perp}$ is thick in $A$-Mod ((((this does not need the assumption of Artin algebra)))).
 Proof.
@@
 Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{GP}$ and ${\rm pd}K=n-1$.
-(normal-(((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{GP}$-approximation of $M$.
+((((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{GP}$-approximation of $M$.
 The following is another theorem of Šaroch - Št́ovíček type (compare Theorem 2.5 and [28, Theorem 4.11]).
@@
 (2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{A}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$.
-(normal-(((If $n=0$, then $K=0.)$
+((((If $n=0$, then $K=0.)$
 (3)  $\operatorname{Ext}_{A}^{1}(M,C)=0$ for every module $C\in\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}.$
@@
 Since $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B},$ it follows that it is also hereditary (cf.
-[27, Lemma 6.17]; or Subsection 1.2).
+[29, Lemma 6.17]; or Subsection 1.2).
 ∎
@@
 Corollary 6.4.
-Let $R$ be a ring. Then the following are equivalent:normal-:::
+Let $R$ be a ring. Then the following are equivalent::::
 (1)  $\mathcal{PGF}=\mathcal{GP}.$
@@
 7.1. Application 4: Exact model structures on $\mathcal{GF}_{n}$
-The following result is due to R. EI Maaouy [22]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [28, Theorem 4.11].
+The following result is due to R. El Maaouy [22]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [28, Theorem 4.11].
 Theorem 7.1.
-([22, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent:normal-:::
+([22, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent::::
 (1)  $M\in\mathcal{GF}_{n}.$
@@
 We stress that the proof of [22, Theorem 3.4] completely holds for Theorem 7.1, thus we will not include a proof.
-When $n=0$, the following result is [28, Theorem 4.9].
-Theorem 7.2.
+When $n=0$, the following result is [28, Corollary 4.12].
+Corollary 7.2.
 ([22, Theorems A, B])  Let $R$ be a ring and $n$ a non-negative integer. Then
@@
 $(\mathcal{P}_{m},\ \mathcal{P}_{m}^{\perp}),\ \ \ \ (\mathcal{F}_{m},\ \mathcal{F}_{m}^{\perp}),\ \ \ \ (\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{PGF}^{\perp}),\ \ \ \ (\mathcal{GF}_{m},\ \mathcal{F}_{m}^{\perp}\cap\mathcal{PGF}^{\perp})$
-where $0\leq m\leq n$ (see Lemma 2.2, Lemma 7.3, Corollary 2.6, and Theorem 7.2). Applying Theorem 5.1 to $\mathcal{B}=\mathcal{GF}_{n}$, one gets the assertions (1) - (4) in the following corollary.
+where $0\leq m\leq n$ (see Lemma 2.2, Lemma 7.3, Corollary 2.6, and Corollary 7.2).
+Applying Theorem 5.1 to $\mathcal{B}=\mathcal{GF}_{n}$, one gets the assertions (1) - (4) in the following corollary.
 Corollary 7.7.

### v4

@@
 As a byproduct, $\mathcal{PGF}=\mathcal{GP}$ for a ring $R$ if and only if $\mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n}=\mathcal{P}_{n}$ for some $n$.
-Keywords: projectively coresolved Gorenstein flat (PGF) module; module of finite PGF (Gorenstein projective, Gorenstein flat) dimension; (complete, hereditary) cotorsion pair; abelian (exact) model structure; weakly idempotent exact category; finitistic dimension
+Keywords: projectively coresolved Gorenstein flat (PGF) module; module of finite PGF (Gorenstein projective, Gorenstein flat) dimension; (complete, hereditary) cotorsion pair; abelian (exact) model structure; weakly idempotent complete exact category; finitistic dimension
 2020 Mathematics Subject Classification. Primary 16E30, 18N40; Secondary 16E10, 16E65, 16G50
@@
 This kind of PGF modules enjoys pleasant properties.
-It is not clear whether a Gorenstein projective module is Gorenstein flat.
-But PGF modules provide a common refinement of Gorenstein projective modules and Gorenstein flat modules: by definition they are Gorenstein flat, and they also prove to be Gorenstein projective ([28, Theorem 4.4]).
+It is not clear whether a Gorenstein projective module is Gorenstein flat; but PGF modules provide a common refinement of Gorenstein projective modules and Gorenstein flat modules: by definition they are Gorenstein flat, and they also prove to be Gorenstein projective ([32, Theorem 4.4]).
 The full subcategory $\mathcal{PGF}$ of $R$-Mod consisting of the PGF modules is a Frobenius category with $\mathcal{P}$, the full subcategory of projective $R$-modules, as the class of projective-injective objects, and the stable triangulated category $\mathcal{PGF}/\mathcal{P}$ is the homotopy category of infinite different abelian or exact model structures, in the sense of D.
@@
 Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra.
-Šaroch and Št́ovíček [28, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$.
+Šaroch and Št́ovíček [32, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$.
 By a result of A.
-Iacob [20, Proposition 9] (see also [22, Proposition 4.11]), if $R$ is a Gorenstein ring (i.e., $R$ is a left and right noetherian ring and the injective dimensions of ${}_{R}R$ and $R_{R}$ are finite), then $\mathcal{PGF}=\mathcal{GP}$.
-Thus, in this case, the cotorsion pair $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ reads as $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, which has been given by Hovey [19, Theorem 8.3].
+Iacob [22, Proposition 9] (see also [25, Proposition 4.11]), if $R$ is a Gorenstein ring (i.e., $R$ is a left and right noetherian ring and the injective dimensions of ${}_{R}R$ and $R_{R}$ are finite), then $\mathcal{PGF}=\mathcal{GP}$.
+Thus, in this case, the cotorsion pair $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ reads as $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, which has been given by Hovey [21, Theorem 8.3].
 By the Hovey correspondence, all these three kinds of cotorsion pairs induce abelian model structures in $R$-Mod.
@@
 $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})\ \ \ \ \mbox{ and}\ \ \ \ (\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$
-are complete and hereditary cotorsion pairs in $R$-Mod (see [11, Theorem 7.4.6], and [23, Theorem 3.4(2)], respectively), where $\mathcal{P}_{n}$ (respectively, $\mathcal{F}_{n})$ is the class of $R$-modules of projective (respectively, flat) dimension $\leq n$.
+are complete and hereditary cotorsion pairs in $R$-Mod (see [12, Theorem 7.4.6], and [26, Theorem 3.4(2)], respectively), where $\mathcal{P}_{n}$ (respectively, $\mathcal{F}_{n})$ is the class of $R$-modules of projective (respectively, flat) dimension $\leq n$.
 The first aim of this paper is to show that these phenomena are quite common. For each non-negative integer $n$, we prove that
@@
 R.
-El Maaouy [22, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
+El Maaouy [25, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
 $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
@@
 $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$
-Enochs, Jenda [11, Theorem 7.4.6]
+Enochs, Jenda [12, Theorem 7.4.6]
 If ${\rm Fpd}<\infty$, then
@@
 $(\mathcal{FL},\ \mathcal{EC})$
-Enochs, Jenda [11, Proposition 7.4.3]
+Enochs, Jenda [12, Proposition 7.4.3]
 $(\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$
-Mao, Ding [23, Theorem 3.4(2)]
+Mao, Ding [26, Theorem 3.4(2)]
 If ${\rm Fpd}<\infty$, then
@@
 $(\mathcal{PGF},\ R\mbox{-}{\rm Mod},\ \mathcal{PGF}^{\perp})$
-Šaroch, Št́ovíček [28, Theorem 4.9]
+Šaroch, Št́ovíček [32, Theorem 4.9]
 $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
@@
 Šaroch, Št́ovíček
-[28, Theorem 4.11, Corollary 4.12]
+[32, Theorem 4.11, Corollary 4.12]
 $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$
 $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$
-Maaouy [22, Theorem A]
+Maaouy [25, Theorem A]
 If ${\rm Fpd}<\infty$, then
@@
 Denote by $\mathcal{PGF}$ the subcategory of PGF modules.
-By definition $\mathcal{PGF}\subseteq\mathcal{GF}.$ It is proved in [28, Theorem 4.4] that $\mathcal{PGF}\subseteq\mathcal{GP}.$ However, it is not clear whether a Gorenstein projective module is PGF.
+By definition $\mathcal{PGF}\subseteq\mathcal{GF}.$ It is proved in [32, Theorem 4.4] that $\mathcal{PGF}\subseteq\mathcal{GP}.$ However, it is not clear whether a Gorenstein projective module is PGF.
 In fact, $\mathcal{PGF}=\mathcal{GP}$ if and only if $\mathcal{GP}\subseteq\mathcal{GF}$.
@@
 A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$, where $\mathcal{S}$ is a set of modules (or, equivalently, cogenerated by a single module).
-Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([12, Theorem 10]; also [16, Theorem 6.11(b)]: but in [16] “cogenerated” is called “generated”).
-This result has been generalized to Grothendieck category with enough projective objects in [19, Theorem 2.4].
+Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([13, Theorem 10]; also [18, Theorem 6.11(b)]: but in [18] “cogenerated” is called “generated”).
+This result has been generalized to Grothendieck category with enough projective objects in [21, Theorem 2.4].
 A cotorsion pair $(\mathcal{C},\mathcal{F})$ is hereditary, if $\mathcal{C}$ is closed under the kernel of deflations, and $\mathcal{F}$ is closed under the cokernel of inflations.
 For a cotorsion pair $(\mathcal{C},\mathcal{F})$ in an abelian category $\mathcal{A}$ with enough projective objects and injective objects, these two conditions of the heredity are equivalent; and they are equivalent to ${{\rm Ext}}^{2}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$ and $F\in\mathcal{F};$ and also to ${{\rm Ext}}^{i}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$, $F\in\mathcal{F}$, and $i\geq 2$.
-See [13, Proposition 1.2.10].
+See [14, Proposition 1.2.10].
 This result holds also for a complete cotorsion pair in a weakly idempotent complete exact category.
-See [29, Lemma 6.17] (also [27]).
+See [33, Lemma 6.17] (also [31]).
 1.3. Model structures
@@
 It is not necessarily a triangulated category, in general.
-A model structure on an exact category is exact ([14, 3.1]), if cofibrations are exactly inflations with cofibrant cokernel, and fibrations are exactly deflations with fibrant kernel.
+A model structure is trivial if each object is a trivial object.
+The homotopy category of a model structure is zero if and only if it is a trivial model structure ([20, Theorem 1.2.10(iv)]).
+We are only interested in model structures which are not trivial.
+A model structure on an exact category is exact ([15, 3.1]), if cofibrations are exactly inflations with cofibrant cokernel, and fibrations are exactly deflations with fibrant kernel.
 In this case, trivial cofibrations are exactly inflations with trivially cofibrant cokernel, and trivial fibrations are exactly deflations with trivially fibrant kernel.
-If $\mathcal{A}$ is an abelian category, then an exact model structure on $\mathcal{A}$ is exactly an abelian model structure in [19].
-An exact model structure is projective if each object is fibrant, or equivalently, each trivially cofibrant object is projective ([14, 4.5, 4.6]).
-An exact model structure is trivial if each object is a trivial object.
-The homotopy category of a model structure is zero if and only if it is a trivial model structure ([18, Theorem 1.2.10(iv)]).
-We are only interested in exact model structures which are not trivial.
+If $\mathcal{A}$ is an abelian category, then an exact model structure on $\mathcal{A}$ is precisely an abelian model structure in [21].
+An exact model structure is projective if each object is fibrant, or equivalently, each trivially cofibrant object is projective ([15, 4.5, 4.6]).
 A Hovey triple ([21], [15]) in an exact category $\mathcal{A}$ is a triple  $(\mathcal{C},\mathcal{F},\mathcal{W})$  of classes of objects such that   $\mathcal{W}$ is thick in $\mathcal{A}$ (i.e., $\mathcal{W}$ is closed under direct summands, and if two out of three terms in an admissible exact sequence are in $\mathcal{W}$, then so is the third one); and that both $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})$ and  $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})$   are complete cotorsion pairs.
@@
 This has been extended to weakly idempotent complete exact category in J.
-Gillespie [14, Theorem 3.3] (see also [29, Theorem 6.9]).
+Gillespie [15, Theorem 3.3] (see also [33, Theorem 6.9]).
 Namely, there is a one-to-one correspondence between exact model structures and the Hovey triples in weakly idempotent complete exact category $\mathcal{A}$, given by
@@
 An advantage of a hereditary Hovey triple is that the homotopy category of the corresponding exact model structure is just the stable category of a Frobenius category, and hence triangulated.
-See Lemma 1.3 below.
-Gillespie [15, Theorem 1.1] has constructed all the hereditary Hovey triples in an abelian category via two compatible complete cotorsion pairs.
+See Theorem 1.3 below.
+Gillespie [16, Theorem 1.1] has constructed all the hereditary Hovey triples in an abelian category via two compatible hereditary complete cotorsion pairs.
+Furthermore, H.
+Nakaoka and Y.
+Palu [NP] introduce an extriangulated category, which is a common generalization of an exact category and a triangulated category, and generalize the Hovey correspondence to this setting up; moreover, remarkably, the homotopy category is proved to be triangulated, even if the two compatible complete cotorsion pairs are not necessarily hereditary.
+See [27, Theorem 6.20], and [17, Theorem 6.34].
 1.4. Homotopy categories of exact model structures
@@
 Then
-$(1)$   ([29, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
+$(1)$   ([33, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
 $(2)$  Put  $\mathcal{S}=\{f\in{\rm Mor}(\mathcal{A})\ |\ f\ \text{is admissible},\ {\rm Ker}f\in\mathcal{W},\ {\rm Coker}f\in\mathcal{W}\}.$ Then
@@
 This follows from Lemma 1.1 and the definition of the homotopy category. ∎
-For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([29, 6.15]; compare [18, 6.2], or, cofibration sequences and fibration sequences in [24]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5].
-The following nice result can be explicitly founded in Št́ovíček [29, Theorem 6.21].
-See also Gillespie [14, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
+For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([33, 6.15]; compare [20, 6.2], or, cofibration sequences and fibration sequences in [28]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5].
+The following nice result can be explicitly founded in Št́ovíček [33, Theorem 6.21].
+See also Gillespie [15, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
 Theorem 1.3.
@@
 This is true, since by Theorem 1.3, the triangulation is determined by the cofiber sequences, which are by definition uniquely determined by conflations in $\mathcal{A}$ and by $\mathcal{W}$.
-See [29, Definition 6.15].
+See [33, Definition 6.15].
 ∎
@@
 Lemma 2.1.
-([28, Theorem 4.9])  Let $R$ be a ring.
+([32, Theorem 4.9])  Let $R$ be a ring.
 Then $\mathcal{PGF}^{\perp}$ is thick; and $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair with kernel $\mathcal{P}$, and cogenerated by a set.
@@
 Lemma 2.2.
-([11, Theorem 7.4.6])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete and hereditary cotorsion pair, and cogenerated by a set.
-The PGF version of Auslander - Buchweitz Theorem ([2, Theorem 1.1]; see also [17, Theorem 2.10]) can be found in G. Dalezios and I. Emmanouil [8].
+([12, Theorem 7.4.6])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete and hereditary cotorsion pair, and cogenerated by a set.
+The PGF version of Auslander - Buchweitz Theorem ([2, Theorem 1.1]; see also [19, Theorem 2.10]) can be found in G. Dalezios and I. Emmanouil [8].
 Lemma 2.3.
@@
 ${\rm(iii)}$   ${\rm PGFd}M_{3}\leq{\rm max}\{{\rm PGFd}M_{1}+1,{\rm PGFd}M_{2}\}$.
-The following result describes the modules of finite PFG dimension $\leq n$. It is a theorem of Šaroch - Št́ovíček type: [28, Theorem 4.11] is the similar result, for Gorenstein flat modules.
+The following result describes the modules of finite PGF dimension $\leq n$. It is a theorem of Šaroch - Št́ovíček type: [32, Theorem 4.11] is the similar result, for Gorenstein flat modules.
 Theorem 2.5.
@@
 ∎
-When $n=0$, the following result coincides with [28, Theorem 4.9].
+When $n=0$, the following result coincides with [32, Theorem 4.9].
 Corollary 2.6.
@@
 To prove that it is complete, it suffices to show that it is cogenerated by a set.
-By [28, Theorem 4.9] (cf.
+By [32, Theorem 4.9] (cf.
 Lemma 2.1), the cotorsion pair  $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is cogenerated by a set, say $T$.
@@
 ∎
-Auslander - Buchweitz Theorem is fundamental in Gorenstein homological algebra. See [2, Theorem 1.1], or more explicitly in [17, Theorem 2.10].
+Auslander - Buchweitz Theorem is fundamental in Gorenstein homological algebra. See [2, Theorem 1.1], or more explicitly in [19, Theorem 2.10].
 Lemma 3.2.
@@
 ((((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{GP}$-approximation of $M$.
-The following is another theorem of Šaroch - Št́ovíček type (compare Theorem 2.5 and [28, Theorem 4.11]).
+The following is another theorem of Šaroch - Št́ovíček type (compare Theorem 2.5 and [32, Theorem 4.11]).
 Theorem 3.3.
@@
 Conversely, if $M\in\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}$, then by Lemma 3.2 there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$.
-Since $K\in\mathcal{P}_{n-1}\subseteq\mathcal{GP}^{\perp}$ and $M\in\mathcal{GP}^{\perp}$, $G\in\mathcal{GP}^{\perp}$, and hence $G\in\mathcal{GP}\cap\mathcal{GP}^{\perp}=\mathcal{P}_{n}.$ Thus $M\in\mathcal{P}_{n}$.
+Since $K\in\mathcal{P}_{n-1}\subseteq\mathcal{GP}^{\perp}$ and $M\in\mathcal{GP}^{\perp}$, $G\in\mathcal{GP}^{\perp}$, and hence $G\in\mathcal{GP}\cap\mathcal{GP}^{\perp}=\mathcal{P}.$ Thus $M\in\mathcal{P}_{n}$.
 $(1)\Longrightarrow(2)$:  Assume that $M\in\mathcal{GP}_{n}.$ The case for $n=0$ is trivial.
@@
 $(4)\Longrightarrow(1)$: Assume that  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ is an exact sequence with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{GP}$.
-Since $L\in\mathcal{P}_{n}\subseteq\mathcal{GP}_{n}$ and $N\in\mathcal{GP}\subseteq\mathcal{GP}_{n}$, it follows from [17, Theorem 2.20] that $M\in\mathcal{GP}_{n}$.
+Since $L\in\mathcal{P}_{n}\subseteq\mathcal{GP}_{n}$ and $N\in\mathcal{GP}\subseteq\mathcal{GP}_{n}$, it follows from [19, Theorem 2.20] that $M\in\mathcal{GP}_{n}$.
 $(2)\Longrightarrow(3):$  Assume that $0\longrightarrow K\stackrel{{\scriptstyle f}}{{\longrightarrow}}G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n}$ is an exact sequence which remains exact after applying $\operatorname{Hom}_{A}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$.
@@
 It is hereditary, since $\mathcal{GP}_{n}$ is closed under taking kernels of epimorphisms (cf.
-[17, Theorem 2.20]); or equivalently $\mathcal{P}_{n}^{\perp}$, and hence $\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$, is closed under taking cokernels of monomorphisms.
+[19, Theorem 2.20]); or equivalently $\mathcal{P}_{n}^{\perp}$, and hence $\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$, is closed under taking cokernels of monomorphisms.
 To prove that it is complete, it suffices to show that for $M\in A\mbox{-}{\rm Mod}$, there is an exact sequence
@@
 We do not need to consider the corresponding results for $\mathcal{GP}_{n}$ for Gorenstein ring $R$, by using the cortorsion pair  $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, since in this case one has $\mathcal{GP}=\mathcal{PGF}$ (cf.
-[20, Proposition 9]).
+[22, Proposition 9]).
 4. Several finitistic dimensions
@@
 For any non-negative integer $n$ and any ring $R$, we have seen in Corollary 2.6 that $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair in $R\mbox{-}{\rm Mod}$, and that $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple in $R\mbox{-}{\rm Mod}.$ Also, for Artin algebra $A$, we have seen in Theorem 3.4 that $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A\mbox{-}{\rm Mod}$, and that $(\mathcal{GP}_{n},\ \mathcal{P}_{n}^{\perp},\ \mathcal{GP}^{\perp})$ is a hereditary Hovey triple in $A\mbox{-}{\rm Mod}.$
-For any ring $R$, Maaouy [22, Theorem A] proves that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R$-Mod.
-It seems to be natural to ask whether $\mathcal{PGF}^{<\infty}$, $\mathcal{GF}^{<\infty}$ and $\mathcal{GF}^{<\infty}$ could be the left parts of cotorsion pairs in $R$-Mod.
+For any ring $R$, Maaouy [25, Theorem A] proves that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R$-Mod.
+It seems to be natural to ask whether $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$ and $\mathcal{GF}^{<\infty}$ could be the left parts of cotorsion pairs in $R$-Mod.
 This problem reasonably leads to several finitistic dimensions.
@@
 ${\rm Ffd}={\rm sup}\{{\rm fd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm fd}M<\infty\}.$
-The (left) finitistic Gorenstein projective dimension ${\rm FGpd}={\rm FGpd}(R)$ ([17, 2.27]) is defined as
+The (left) finitistic Gorenstein projective dimension ${\rm FGpd}={\rm FGpd}(R)$ ([19, 2.27]) is defined as
 ${\rm FGpd}={\rm sup}\{{\rm Gpd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm Gpd}M<\infty\}.$
-The (left) finitistic Gorenstein flat dimension FGfd = FGfd$(R)$ ([17, 3.24]) is defined as
+The (left) finitistic Gorenstein flat dimension FGfd = FGfd$(R)$ ([19, 3.24]) is defined as
 ${\rm FGfd}={\rm sup}\{{\rm Gfd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm Gfd}M<\infty\}.$
@@
 ${\rm FPGFd}={\rm sup}\{{\rm PGFd}M\ |\ M\in R\mbox{-}{\rm Mod},\ {\rm PGFd}M<\infty\}.$
-Using an argument of H. Holm, one can prove the following analogue of ([17, Theorem 2.28]).
 Proposition 4.1.
 Let $R$ be a ring. Then
-(1)  ${\rm FPGFd}={\rm FGpd}={\rm Fpd}$.
-(2)  If $R$ is right coherent and ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$.
-(3)  If $R$ is right coherent, then ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
+(1)  One has ${\rm FPGFd}={\rm FGpd}={\rm Fpd}$.
+(2)  If ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$.
+(3)  One has ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
 Proof.
-(1)  [17, Theorem 2.28] claims that ${\rm FGpd}={\rm Fpd}$. Thus, it suffices to prove ${\rm FPGFd}={\rm Fpd}$.
+(1)  [19, Theorem 2.28] claims that ${\rm FGpd}={\rm Fpd}$. Thus, it suffices to prove ${\rm FPGFd}={\rm Fpd}$.
 Since $\mathcal{P}\subseteq\mathcal{PGF}\subseteq\mathcal{GP}$, it follows that ${\rm Gpd}M\leq{\rm PGFd}M\leq{\rm pd}M$ for any $R$-module.
-If ${\rm pd}M<\infty$, then ${\rm Gpd}M={\rm pd}M$ (see [17, Proposition 2.27]), and hence ${\rm PGFd}M={\rm pd}M$ (see also [12, Corollary 3.7(i)]). It follows that ${\rm Fpd}\leq{\rm FPGFd}.$
+If ${\rm pd}M<\infty$, then ${\rm Gpd}M={\rm pd}M$ (see [19, Proposition 2.27]), and hence ${\rm PGFd}M={\rm pd}M$ (see also [13, Corollary 3.7(i)]). It follows that ${\rm Fpd}\leq{\rm FPGFd}.$
 If $0<{\rm PGFd}M=t<\infty$, then by [8, Proposition 3.1] (cf.
@@
 Thus ${\rm pd}L=m$.
-(2)   Assume that ${\rm fd}M=m<\infty$.
-Since $\mathcal{F}\subseteq\mathcal{GF}$, it follows that ${\rm Gfd}M\leq{\rm fd}M=m$.
-To prove ${\rm Gfd}M=m$, by [17, Theorem 3.14] it suffices to find an injective right $R$-module $I$ such that ${\rm Tor}^{R}_{m}(I,M)\neq 0$.
-Since ${\rm fd}M=m$, it follows that there is a right $R$-module $N$ such that ${\rm Tor}^{R}_{m}(N,M)\neq 0$.
-Consider an exact sequence $0\longrightarrow N\longrightarrow I\longrightarrow C\longrightarrow 0$ such that $I$ is an injective right $R$-module.
-Then by the exact sequence
-${\rm Tor}^{R}_{m+1}(C,M)=0\longrightarrow{\rm Tor}^{R}_{m}(N,M)\longrightarrow{\rm Tor}^{R}_{m}(I,M)\longrightarrow{\rm Tor}^{R}_{m}(C,M)$
-one sees that ${\rm Tor}^{R}_{m}(I,M)\neq 0$. We are done.
-(3)  [17, Theorem 3.24] already claims that if $R$ is right coherent, then ${\rm FGfd}={\rm Ffd}$.
-Since $\mathcal{F}\subseteq\mathcal{PGF}\subseteq\mathcal{GF}$, it follows that ${\rm Gfd}M\leq{\rm PGFd}M\leq{\rm fd}M$ for any $R$-module.
-If ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$, by (2).
-It follows that ${\rm PGFd}M={\rm fd}M$.
-This implies that ${\rm Ffd}\leq{\rm FPGFd}.$ Together with (1) one gets ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
+(2)  According to [10, Proposition 3.8], if $M\in\mathcal{PGF}^{\perp}$ then ${\rm Gfd}M={\rm fd}M$.
+Thus, it suffices to say that ${\mathcal{F}}^{<\infty}\subseteq\mathcal{PGF}^{\perp}$.
+By [32, Theorem 4.4 or Theorem 4.11], ${\mathcal{F}}\subseteq{\rm PGF}^{\perp}$.
+Then ${\mathcal{F}}^{<\infty}\subseteq\mathcal{PGF}^{\perp}$ since $\mathcal{PGF}^{\perp}$ is thick.
+(3)  I. Emmanouil [10, Proposition 2.3] already claims that ${\rm FGfd}={\rm Ffd}$.
+By C. U. Jensen [23, Proposition 6], if ${\rm Fpd}<\infty$ then all flat modules have finite projective dimension. It follows that ${\rm Ffd}\leq\mathcal{\rm Fpd}$.
+Together with (1) one gets ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
+Remark 4.2.
+We thank Ioannis Emmanouil for pointing out an error in the original proof of Proposition 4.1(3), and for pointing out to us the paper [23].
 4.2. Abelian model structures arising from $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$, and $\mathcal{GF}^{<\infty}$
@@
 (2)  If $R$ is an Artin algebra, then  $(\mathcal{GP}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp}\cap\mathcal{GP}^{\perp})=(\mathcal{GP}_{t},\ \mathcal{P}_{t}^{\perp}\cap\mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{GP}^{<\infty},\ (\mathcal{P}^{<\infty})^{\perp},\ \mathcal{GP}^{\perp})=(\mathcal{GP}_{t},\ \mathcal{P}_{t}^{\perp},\ \mathcal{GP}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
-(3)  If $R$ is right coherent, then  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
+(3)  The pair  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
 Proof.
-(1)  By Proposition 4.1, one has ${\rm FGFd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}=t$, thus $\mathcal{PGF}^{<\infty}=\mathcal{PGF}_{t}$ and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  $\mathcal{GP}^{<\infty}=\mathcal{GP}_{t}$  and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  and $\mathcal{GF}^{<\infty}=\mathcal{GF}_{t}$  and $\mathcal{F}^{<\infty}=\mathcal{F}_{t}$.
-Then the assertions follow from Corollary 2.6, Theorem 3.4, and [22, Theorem A], respectively.∎
+By Proposition 4.1(3) one has ${\rm FGFd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}=t$.
+Thus $\mathcal{PGF}^{<\infty}=\mathcal{PGF}_{t}$ and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  $\mathcal{GP}^{<\infty}=\mathcal{GP}_{t}$  and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  and $\mathcal{GF}^{<\infty}=\mathcal{GF}_{t}$  and $\mathcal{F}^{<\infty}=\mathcal{F}_{t}$.
+Then the assertions (1), (2) and (3) follow from Corollary 2.6, Theorem 3.4, and [25, Theorem A], respectively.∎
 5. Induced complete cotorsion pairs in exact categories with applications
@@
 Since $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B},$ it follows that it is also hereditary (cf.
-[29, Lemma 6.17]; or Subsection 1.2).
+[33, Lemma 6.17]; or Subsection 1.2).
 ∎
@@
 Since $X\in\mathcal{P}_{n}$ and $L\in\mathcal{P}_{n}$, it follows that ${\rm pd}G<\infty.$ While $G\in\mathcal{PGF}$, $G$ is Gorenstein projective, and hence $G$ has to be projective (cf.
-[11, Proposition 10.2.3]).
+[12, Proposition 10.2.3]).
 Thus $Z\in\mathcal{P}_{n}$.
@@
 Note that $\mathcal{GP}_{n}$ is closed under extensions, the kernels of epimorphisms, and direct summands (cf.
-[17, Theorem 2.20, Proposition 2.19]).
+[19, Theorem 2.20, Proposition 2.19]).
 In particular, $\mathcal{GP}_{n}$ is a weakly idempotent complete exact category.
@@
 Also note that $\mathcal{GP}_{m}\cap\mathcal{P}^{<\infty}=\mathcal{P}_{m}$ (cf.
-[17, Proposition 2.27]).
+[19, Proposition 2.27]).
 Similar to Corollary 6.1 and Theorem 6.3 one has
@@
 7.1. Application 4: Exact model structures on $\mathcal{GF}_{n}$
-The following result is due to R. El Maaouy [22]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [28, Theorem 4.11].
+The following result is due to R. El Maaouy [25]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [32, Theorem 4.11].
 Theorem 7.1.
-([22, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent::::
+([25, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent::::
 (1)  $M\in\mathcal{GF}_{n}.$
@@
 (4)  There is an exact sequence  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ with $L\in\mathcal{F}_{n}$ and $N\in\mathcal{PGF}$.
-The careful reader will find that the statement (2) is slightly stronger than [22, Theorem 3.4]: here $K\in\mathcal{F}_{n-1}$ if $n\geq 1$, rather than $K\in\mathcal{F}_{n}$ for all $n\geq 0$ as in [22, Theorem 3.4]; also, $X\in\mathcal{F}_{n}^{\perp}$ in (2), rather than $X\in\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp}$ as in [22, Theorem 3.4].
+The careful reader will find that the statement (2) is slightly stronger than [25, Theorem 3.4]: here $K\in\mathcal{F}_{n-1}$ if $n\geq 1$, rather than $K\in\mathcal{F}_{n}$ for all $n\geq 0$ as in [25, Theorem 3.4]; also, $X\in\mathcal{F}_{n}^{\perp}$ in (2), rather than $X\in\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp}$ as in [25, Theorem 3.4].
 In this way, Theorem 7.1 coincides with the corresponding versions for modules in $\mathcal{PGF}_{n}$ as in Theorem 2.5, and for modules $\mathcal{GP}_{n}$ as in Theorem 3.3.
-We stress that the proof of [22, Theorem 3.4] completely holds for Theorem 7.1, thus we will not include a proof.
-When $n=0$, the following result is [28, Corollary 4.12].
+We stress that the proof of [25, Theorem 3.4] completely holds for Theorem 7.1, thus we will not include a proof.
+When $n=0$, the following result is [32, Corollary 4.12].
 Corollary 7.2.
-([22, Theorems A, B])  Let $R$ be a ring and $n$ a non-negative integer. Then
+([25, Theorems A, B])  Let $R$ be a ring and $n$ a non-negative integer. Then
 (1)  The pair  $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair in $R\mbox{-}{\rm Mod}$ with kernel $\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp}$.
@@
 Lemma 7.3.
-([23, Theorem 3.4(2)])  Let $R$ be a ring, and $n$ a non-negative integer.
+([26, Theorem 3.4(2)])  Let $R$ be a ring, and $n$ a non-negative integer.
 Then $(\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, and cogenerated by a set.
@@
 By [8, Proposition 2.5(ii)], ${\rm pd}M={\rm fd}M<\infty$.
-Since $M\in\mathcal{PGF}$, it follows that $M$ is Gorenstein projective ([28, Theorem 4.4]), with finite projective dimension, and hence $M$ is projective (cf.
-[11, Proposition 10.2.3]).
+Since $M\in\mathcal{PGF}$, it follows that $M$ is Gorenstein projective ([32, Theorem 4.4]), with finite projective dimension, and hence $M$ is projective (cf.
+[12, Proposition 10.2.3]).
 An alternative proof: by Theorem 7.1 one has $M\in\mathcal{PGF}\cap\mathcal{F}_{n}=\mathcal{PGF}\cap\mathcal{GF}_{n}\cap\mathcal{PGF}^{\perp}=(\mathcal{PGF}\cap\mathcal{PGF}^{\perp})\cap\mathcal{GF}_{n}=\mathcal{P}\cap\mathcal{GF}_{n}=\mathcal{P}.$
@@
 The corresponding homotopy category is  $(\mathcal{GF}_{m}\cap\mathcal{F}_{m}^{\perp})/(\mathcal{F}_{m}\cap\mathcal{F}_{m}^{\perp})\cong\mathcal{PGF}/\mathcal{P}.$
+Acknowledgement: We thank Ioannis Emmanouil for his useful comments.